### v2

# REACT: Real-time Efficiency and Accuracy Compromise for Tradeoffs in Scene Graph Generation

Abstract
Scene Graph Generation (SGG) is a task that encodes visual relationships between objects in images as graph structures. SGG shows significant promise as a foundational component for downstream tasks, such as reasoning for embodied agents. To enable real-time applications, SGG must address the trade-off between performance and inference speed. However, current methods tend to focus on one of the following: (1) improving relation prediction accuracy, (2) enhancing object detection accuracy, or (3) reducing latency, without aiming to balance all three objectives simultaneously. To address this limitation, we propose a novel architecture, inference method, and relation prediction model. Our proposed solution, the REACT model, achieves the highest inference speed among existing SGG models, improving object detection accuracy without sacrificing relation prediction performance. Compared to state-of-the-art approaches, REACT is 2.7 times faster (with a latency of 23 ms) and improves object detection accuracy by 58.51%. Furthermore, our proposal significantly reduces model size, with an average of 5.5x fewer parameters [^1] Code is available at https://github.com/Maelic/SGG-Benchmark.
1 Introduction
Figure 1: Performance comparison between our approach and previous work, in terms of performance (F1@K metric) and Latency (ms) on the PSG dataset [42]. The points are scaled by the number of parameters for each model (smaller is better).
Scene Graph Generation (SGG) is the task of generating a structured graph representation from visual inputs. Given an image, visual regions corresponding to objects of interest are extracted, and relations between these objects are predicted in the form of $<subject,predicate,object>$ triplets. Aggregated together, these triplets form a directed acyclic graph which can be interpreted as a structured representation of the scene. In addition, subject and object nodes are traditionally associated with their coordinates on the image (i.e., bounding boxes). In recent years, this task has garnered increasing attention thanks to its utility in various downstream applications, such as Visual Question Answering (VQA) [8, 5, 23], Image Captioning [43, 40, 30], or Reasoning of an embodied agent [7, 22, 28]. However, we observe a significant unbalance between the progress in the SGG task itself and its adoption as a foundational component for downstream tasks [2]. One of the reasons is the lack of low-cost approaches that could power applications with real-time constraints. Because SGG is a task that combines Object Detection (OD) and Relation Prediction (RelPred), a good trade-off between accuracy in these two parameters is also required. Nonetheless, current approaches focus on one or two of these parameters without considering the general problem. In this work, we explore the bottlenecks in SGG to find an optimal combination of OD accuracy, RelPred accuracy, and latency (see Figure 1).
We can separate SGG methods into two categories: Two-Stage approaches (TS) and One-Stage approaches (OS). The former uses a combination of an object detector and a relation predictor to model objects and relations in a sequence. The latter uses a single-stage pipeline, to infer relations and object proposals directly from the image features. As a result, one-stage approaches are often more efficient for real-time processing [10]. However, a few challenges remain, especially concerning the accuracy of such approaches in the OD task [20, 9]. In contrast, two-stage approaches train the object detector and relation predictor sequentially, leading to improved performance in object detection.
In this work, we propose to investigate the bottlenecks of the two-stage architecture, in terms of OD accuracy and latency. First, structural dependencies between blocks of the architecture are reviewed as a Decoupled Two-Stage SGG (DTS-SGG). This design permits to maintain performance of the pre-trained detector and allows better control of the information flow between the two stages of the architecture. DTS enables the selection of a smaller set of high-quality objects (i.e., proposals) before passing them to the relation prediction stage. This approach reduces overall latency, as the complexity of relation prediction increases exponentially with the number of input objects. To select an optimal amount of proposals at inference time, we define a new algorithm, the Dynamic Candidate Selection (DCS). DCS uses evaluation results to find the best trade-off between latency and accuracy at test time. Thanks to the DTS architecture, we can switch the dependency on a Faster-RCNN model for any real-time object detectors, such as one of the YOLO family (YOLOV8/9/10 [15, 39, 38]). Lastly, popular architectures in SGG [46] use a combination of different feature sources, such as union pair features [44] or textual features of the subject and object classes [27]. This results in a significant amount of processing to gather and refine each type of feature altogether. By removing the dependence on sub-optimal features, we can reduce the parameter count of the model and lower the latency. By coupling this strategy to the DTS architecture, we propose a new model for SGG, the Real-Time Efficiency and Accuracy Compromise for Tradeoffs in Scene Graph Generation (REACT). The contributions of this work are three-fold:
  1. 1.
    A new architecture for SGG, the Decoupled Two-Stage (DTS). Thanks to this new architecture, real-time object detectors can be coupled with state-of-the-art two-stage RelPred components, and lead to a gain of latency of up to 10x compared to baseline models.
  2. 2.
    A new inference method, Dynamic Candidate Selection (DCS), which finds an optimal number of proposals to select for the relation prediction stage, to further improve latency without compromising accuracy.
  3. 3.
    A new model, REACT, that extends previous work and our DTS architecture. REACT achieves reduced latency to just 23ms (44 FPS) while enhancing performance on standard metrics (F1@K).
Our findings demonstrate the inefficiency of current implementations of SGG models and even highlight a few drawbacks in the design of prior approaches, which can be taken into account for future work. We also want to emphasize that our method can considerably reduce the time and resources needed for training, which is a step toward more sustainable SGG implementations [33].
2 Related Work
Two-stage approaches rely on a Faster-RCNN detector with a ResNeXt-101 backbone for object detection and feature extraction [41]. Classical approaches include Neural-Motifs [44], VCTree [34], Transformer [35], GPS-Net [26] or PE-NET [46]. Faster-RCNN is not optimal for real-time constraints because of the size and complexity of the backbone. In recent work, Jin et al. [11] introduced an approach for real-time SGG based on contextual information. In contrast to other approaches, this work does not use visual features to infer relations. Instead, it uses the object’s bounding box coordinates predicted by a YOLOV5 [14] model to learn the correlation between pairs of coordinates and predicate classes. However, the reliance on spatial features only will neglect fine-grained relations such as the difference between $eating$ or $drinking$. Another work [13] proposed to achieve real-time SGG by leveraging a Relation-aware YOLO structure (RYOLO). The approach involves predicting the coordinates of relations by adding new anchors on the feature maps of YOLO. Then, matching is done with spatially proximate objects detected. However, such an approach leads to poor performance in RelPred, with only 3.1% in the mR@100 metric. This can be due to the absence of contextual information from the subject and object during relation prediction [44]. In contrast, we propose to use the features extracted by a YOLO model to generate representations for the subject and object nodes, making the first stage of the architecture more efficient without compromising performance on the second stage.
Recently, the paradigm of SGG has been tackled as a one-stage architecture, learning to represent both objects and relations at the same time. For instance, Sparse-RCNN [36] employs a strategy of triplet querying to generate object proposals and relations. Once visual features have been extracted from a CNN backbone, boxes and relations are decoded altogether using a cascade-RCNN scheme. RelationTransformer (RelTr) [3] and SGTR [21] are two approaches based on the DETR architecture [1]. Both approaches generate a sparse set of relations between $<subject,object>$ pairs, in contrast to traditional two-stage approaches which will compute probabilities for all possible relations (dense predictions). Inspired by sparse approaches in one-stage SGG, we aim to reduce the computational bottleneck of two-stage SGG by selecting only relevant object pairs for the relation prediction stage. Lately, Im et al. introduced EGTR [10], a new one-stage approach based on a Deformable-DETR architecture, more efficient than classical DETR for the OD task, paired with a graph relation matching process for RelPred. Authors of RelTr and EGTR have compared their work with two-stage approaches such as Motifs [44] and VCTree [34] in terms of latency. The comparison shows a clear disadvantage for any two-stage approaches. This may be biased by the use of Faster-RCNN for OD in two-stage approaches. We strongly believe that lower latency and better accuracy can be attained by modifying two-stage approaches to incorporate real-time object detectors such as YOLO in place of Faster-RCNN.
(a) Stage 1 - Two-Stage Faster-RCNN
(b) Stage 2 - Prototype Learning (PE-NET)
(c) Stage 1 - DTS YOLO (ours)
(d) Stage 2 - Efficient Prototype Learning (ours)
Figure 2: Top: the PE-NET architecture [46]. Bottom: our proposed REACT model with (left) Decoupled Two-Stage (DTS) + YOLO for efficient object detection and feature extraction ; (right) our proposed Efficient Prototype Learning model. [- - -] represent modified or added components, $\bigotimes$ denotes element-wise concatenation.
3 Investigating Bottlenecks in SGG
In Figure 2 (top row) we display as current the PE-NET model [46]: a state-of-the-art two-stage SGG approach. The feature extraction (stage 1 Figure 2(a)) as well as the graph prediction (module 4, Figure 2(b)) is common to all two-stage architectures [34, 35, 6, 19, 44, 26]. In Stage 1, visual features are extracted by a ResNeXt-101 backbone (Figure 2(a) (a)), and a Feature Pyramid Network (FPN) [25] is used to refine the features at different scales (Figure 2(a) (b)). Then, a Region Proposal Network (RPN) [32] generates proposals (Figure 2(a) (c)) which are used to align three different types of features with corresponding image regions: union features, spatial features, and visual features (Figure 2(a) (d), (e), and (f)). Union and spatial features are aggregated for every possible pair of proposals and will serve as the initial edge representation in Stage 2 (right) which is the relation prediction stage (Figure 2(b)). A classifier extracts the class labels of each proposal (Figure 2(b) (h)) and retrieves corresponding linguistic features from pre-trained models (GloVe word embeddings [31]). Then, the visual features are upsampled (Figure 2(b) (i)) and combined with the textual features to form the node representation through semantic fusion (Figure 2(b) (j)). Then, prototype learning takes place (Figure 2(b) (k)). Finally, object labels and predicate labels are decoded to form the final graph representation (Figure 2(b) (l), (m), and (n)). This architecture is complex and may not be efficient for retaining the OD performance of the backbone detector and providing low latency for real-time applications. In the following, we present in more detail the bottlenecks related to the Object Detection Accuracy, Proposal Candidate Selection, and the latency of the model.
3.1 Object Detection Accuracy
As we can see in Figure 2(b), the strategy employed in the two-stage SGG is to decode object labels twice during the relation prediction stage: one time for extracting text features (Figure 2(b) (h)) and one time to compute the final predictions (Figure 2(b) (m)). The predictions obtained from this last step depend on each model’s context learning strategy [34, 44, 26, 35]. This architecture design is (1) redundant as similar operations are performed twice and (2) can introduce significant differences in object detection accuracy in contrast to the original Faster-RCNN implementation.
In Table 1, we draw a comparison of the performance of the Faster-RCNN model equipped with a ResNeXt-101 feature extractor on the PSG dataset with different relation prediction heads. We can observe that the performance in OD drops significantly, from 1 point for Motifs-TDE [44, 35] to almost 13 points in mAP@50 for GPS-NET [26] after the relation prediction stage.
Table 1: Performance on Object Detection of Faster-RCNN before and after the relation prediction stage with different SGG models on the PSG dataset [42].
Backbone Relation Head mAP@50
Faster-RCNN - 36.4
Neural-Motifs [44] 35.9
PE-NET [46] 35.4
Transformer [35] 34.6
VC-Tree [34] 32.58
GPS-NET [26] 31.55
These results confirm a true dependency between both stages of the pipeline, which seems to hinder the performance of models in the OD task. Using the Faster-RCNN detector for feature extraction and object detection may not be efficient for low latency and high accuracy. In fact, real-time detectors such as YOLOV8 are today largely outperforming Faster-RCNN [15]. However, simply replacing Faster-RCNN with YOLO in the architecture of Two-Stage SGG is not possible because YOLO does not possess an RPN or ROI Align component.
3.2 Proposal Candidate Selection
Subsequently, performing Non-Maximum Suppression (NMS) (Figure 2(b) (n)) after the context learning is inefficient because a lot of false positive or irrelevant proposals will be selected as node candidates during features refinement and context learning. To alleviate the lack of NMS before stage 1, a fixed set of the top proposals (i.e. 80 or 100) are traditionally considered valid node candidates for stage 2. The idea here is that by sampling a large number of proposals per image, it is easier to find valid pairs and relations. However, this considerably increases the computational complexity required to generate the graph, which is a bottleneck for real-time applications. Having a large number of proposals as candidates can make the final graph less informative, as some important objects can be left out if they do not possess at least one high confidence relation. To alleviate this issue, prior works [44, 35, 19] compute the final ranking of relations with the following score formula:
$\theta_{rel}=\theta_{obj}*\theta_{pred}*\theta_{subj}$ (1)
[…]
3.3 Latency of the PE-NET model
Lastly, we investigate the bottlenecks in latency for the relation prediction stage (see Figure 2(b)) of the PE-NET model [46]. In PE-NET, node prototypes are formed using a combination of linguistic features and visual features for each independent proposal (see Figure 2(b) (j)). Edge features are computed using the concatenation of the spatial and union features (see Figure 2(b) (g)). During prototype learning, prototype representations of the subject and object are combined and added to the edge features to form the relation prototype. Finally, a gate mechanism is employed to select relevant features from the multi-modal fusion. The use of both union and spatial features for the edge representation represents a consequent amount of overload for the model. In Figure 2(a) we see that another step of ROI Align is performed on the union of subject and object boxes in step (d). However, a lot of visual information is already present in the object and subject features (see Figure 2(a) (f)) which will be fused with edge features during prototype learning. In the original work [46], this design is not justified. In addition, during edge and node feature refinement, all features are traditionally upsampled to a fixed size of 4096 (see Figure 2(b) (g) and Figure 2(b) (i)). However, larger feature sizes can significantly augment the latency of the model. Therefore, we aim to investigate the usage and size of each feature type to find an optimal trade-off between latency and performance.
4 Toward Real-Time SGG
Based on observations in Section 3, we propose a novel architecture for real-time SGG, which is depicted in Figure 2(c) and Figure 2(d). We apply the Decoupled Two-Stage architecture and replace the Faster-RCNN detector of PE-NET with a YOLO backbone, as shown in Figure 2(c). We improve the Prototype Learning of PE-NET by removing superfluous components, which results in Efficient Prototype Learning. We name the combination of DTS and Efficient Prototype Learning as the REACT model. We also propose Dynamic Candidate Selection (DCS) to further reduce the latency of the REACT model at inference time. In the following, we detail these components.
4.1 Decoupled Two-Stage SGG
To maintain the object detector accuracy during relation prediction training, we propose to freeze the regression and the classification head of the object detector and perform NMS before the relation prediction stage. With this strategy, the object classifier in the relation prediction stage is also removed and class prediction from the object detector is used for the final prediction. The objective of the relation prediction stage becomes then to predict only the pairs and associate predicates, in contrast to also predicting the class labels of objects. The original loss used during relation prediction: $\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{rel}}+\mathcal{L}_{\text{obj}}$ becomes: $\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{rel}}$. This strategy significantly lowers the complexity and thus the computational load of the model. In addition, this approach helps to fairly evaluate the context learning of SGG models, independently of the object detector used. With our Decoupled Two-Stage architecture, any object detector can be used in place of Faster-RCNN, without impacting stage 2 of the architecture. For instance, one could consider replacing Faster-RCNN with a real-time object detector such as YOLO [11]. Recent versions of YOLO, such as YOLOV8 [15], YOLOV9 [39] or YOLOV10 [38] are very efficient for real-time object detection. To take advantage of YOLO versions, we leveraged the introduced Decoupled Two-Stage strategy and replaced the ResNeXt-101 backbone with YOLO’s CSPNet-based backbones [15, 39] (see Figure 2(c) (a)). P5-FPN of Faster-RCNN was also replaced by the YOLOV8/9/10 PAN-FPN Necks (see Figure 2(c) (b)). Finally, we removed the RPN and detection head of Faster-RCNN and replaced them with the YOLOV8/9/10 Detection Head (see Figure 2(c) (c)). Unlike Faster-RCNN, YOLO-based architectures are not designed for feature extraction because they do not implement any ROI Align module. To address this issue, we combined the feature maps extracted from the P3, P4, and P5 layers of YOLO and added one 3x3 Conv block to reduce the feature sizes before ROI Align (see Figure 2(c) (f)). The ROI Align module from Faster-RCNN has been modified to align features with corresponding bounding boxes at only three different scales (instead of five in the baseline implementation, see Figure 2(c) (f)). Thanks to our DTS strategy, switching from one YOLO version to another does not impact the SGG architecture.
4.2 Dynamic Candidate Selection
Here, we introduce the Dynamic Candidate Selection (DCS) method to select an optimal number of proposal candidates at inference time. During training, the relation prediction stage takes as input a maximum amount of proposals $\theta$ ($\theta=100$). During evaluation, we uniformly sample the number of proposals $k$, starting from 0. Using the generated slope for each metric, we can compute the optimal Dynamic Candidate Selection threshold as follows:
$x_{\text{opt}}=\min\left\{x\mid\left|f^{\prime}(x)\right|<\epsilon\right\}$ (2)
During inference, $x_{\text{opt}}$ is applied to control the maximum number of proposals to use as input for the relation prediction stage, successfully lowering the complexity of the task.
4.3 Efficient Prototype Learning
In PE-NET, node, and edge features are upsampled 8x and 4x respectively to a fixed size of 4096 by a feed-forward layer (see steps (g) and (h) in Figure 2(b)). We believe this is unnecessary and that the model will perform similarly without upsampling. We then removed this layer and kept the original feature sizes of 256 and 512, respectively. In the latter stage during prototype learning, we removed the associated downsampling layers. For the PE-NET model, the use of union features for the edge representation seems redundant with the visual features of the subject and object. As a result, we removed the use of union features and dedicated ROI Align step (see Figure 2(a) (d)). By doing so, the edge representation refinement step (see Figure 2(a) (g)) can also be removed and the spatial features fed directly to the context learning (see Figure 2(d) (i)). To further boost accuracy, we proposed to modify the fusion of subject $s$ and object $o$ features during Prototype Learning of the PE-NET model. We replaced the fusion operation: $ReLU(s+o)-(s-o)^{2}$ with a single linear layer and removed the subsequent gate operation.
Table 2: Full results of our experiments with the YOLOV8 and Faster-RCNN models on the test set of the PSG dataset. $\dagger$ represents evaluation with our DCS strategy. Bold and underlined represent the best and second best for each metric. $\downarrow$ and $\uparrow$ represent the improvements of DTS-YOLO against corresponding Two-Stage baselines with Faster-RCNN. See Figure 1 for a visual representation.
B D Relation Head mR@20/50/100 R@20/50/100 F1@K mAP50 Latency Params
CSPDarkNet-53 YOLOV8m-DTS REACT (ours) $\dagger$  18.3 /  20.1 / 20.8 27.5 / 30.9 / 32.2 23.9 53.1 23.0 43.3M
REACT (ours) 18.3 / 20.0 / 20.9 27.6 / 30.9 / 32.3 23.9 53.1 32.5 43.3M
PE-NET $\dagger$ [46] 17.1 / 19.0 / 19.8 28.0 / 31.3 / 32.9 23.2 $\uparrow$9.1 53.1 $\uparrow$17.6 51.7 $\downarrow$338 187M
PE-NET [46] 17.1 / 19.0 / 19.9 28.0 / 31.3 / 33.0 23.2 $\uparrow$9.1 53.1 $\uparrow$17.6 103.0 $\downarrow$261 187M
GPS-NET $\dagger$ [26] 10.9 / 12.6 / 13.5 26.3 / 30.0 / 31.8 17.3 $\uparrow$5.7 53.1 $\uparrow$21.5 27.8 $\downarrow$286 45.3M
GPS-NET [26] 11.1 / 12.7 / 13.7 26.6 / 30.1 / 31.9 17.6 $\uparrow$5.9 53.1 $\uparrow$21.5 41.1 $\downarrow$272 45.3M
Neural-Motifs $\dagger$ [44] 10.1 / 11.6 / 12.5 25.7 / 28.7 / 30.3 16.2 $\uparrow$3.8 53.1 $\uparrow$17.2 22.4 $\downarrow$296 52.9M
Neural-Motifs [44] 10.1 / 11.6 / 12.5 25.7 / 28.7 / 30.4 16.2 $\uparrow$3.8 53.1 $\uparrow$17.2 35.8 $\downarrow$282 52.9M
VCTree $\dagger$ [34] 10.7 / 12.2 / 12.9 22.1 / 25.3 / 26.8 16.1 $\uparrow$3.7 53.1 $\uparrow$20.5 183.2 $\downarrow$336 268.0M
VCTree [34] 10.7 / 12.2 / 13.0 22.1 / 25.3 / 26.8 16.1 $\uparrow$3.8 53.1 $\uparrow$20.5 239.5 $\downarrow$280 268.0M
Transformer $\dagger$ [35] 11.3 / 12.4 / 12.8 24.1 / 27.0 / 28.3 16.6 $\uparrow$5.5 53.1 $\uparrow$18.5 21.1 $\downarrow$275 97.9M
Transformer [35] 11.3 / 12.4 / 13.2 24.1 / 27.1 / 28.9 16.9 $\uparrow$5.7 53.1 $\uparrow$18.5 35.1 $\downarrow$263 97.9M
ResNeXt-101 Faster-RCNN PE-NET [46] 10.7 / 11.7 / 12.0 16.8 / 18.7 / 19.5 14.1 35.5 390.4 426.5M
GPS-NET [26] 8.2 / 8.8 / 9.2 15.7 / 18.1 / 19.2 11.7 31.6 313.9 391.6M
Neural-Motifs [44] 8.9 / 9.6 / 9.8 16.9 / 18.7 / 19.5 12.5 35.9 318.6 369.6M
VCTree [34] 7.9 / 9.1 / 9.6 18.0 / 20.7 / 22.3 12.4 32.6 519.5 365.9M
Transformer [35] 7.6 / 8.2 / 8.5 16.2 / 17.9 / 18.7 11.1 34.6 296.2 328.3M
DETR EGTR [10] 12.0 / 14.5 / 16.6 24.9 / 30.3 / 33.8 19.4 33.6 78.3 42.5M
RelTR [4] 7.9 / 10.0 / 11.7 12.9 / 17.3 / 20.2 12.4 25.7 59.6 63.7M
5 Experiments
Table 3: Full results of our experiments on the test set of IndoorVG. mR@K and R@K represent the average for $k=[20,50,100]$. $\dagger$ represents evaluation with our DCS strategy.
B Relation Head mR@K R@K F1@K mAP50
DTS REACT (ours) 18.0 30.9 22.8 37.2
PE-NET 16.0 29.4 20.7 37.2
GPS-NET 9.9 29.2 14.8 37.2
Neural-Motifs 11.1 30.5 16.2 37.2
VCTree 14.9 18.9 16.7 37.2
Transformer 12.3 28.4 17.2 37.2
TS PE-NET 13.8 27.1 18.1 25.3
GPS-NET 6.7 11.9 8.6 14.8
Neural-Motifs 9.8 15.2 11.9 26.2
VCTree 11.0 14.7 12.6 25.5
Transformer 9.9 16.8 12.5 24.3
OS EGTR [10] 7.1 10.6 8.5 14.7
RelTR [4] 7.7 11.4 9.2 14.1
Datasets. In SGG, the most widely used dataset is Visual Genome [17]. The commonly used split of the data is called VG150 and contains annotations from the top 150 object classes and top 50 predicate classes. However, VG150 suffers from severe biases [42, 29, 8] and inherent issues such as the presence of ambiguous classes (e.g. the classes ”people”, ”men” etc) [42, 24]. Recently, two new datasets have been proposed with higher-quality annotations: the PSG [42] and IndoorVG [28] datasets. Both datasets have been carefully annotated to remove the presence of ambiguous classes or wrong annotations, which makes them a better choice for benchmarking our approach.
Metrics. Following previous work [41, 34], we used the Recall@K (R@K) and meanRecall@K (mR@K) metrics to measure the performance of models. For both metrics, we evaluate the recall on the top $K$ ($k=[20,50,100]$) relations predicted, ranked by confidence. The Recall@K evaluates the overall performance of a model on the selected dataset whereas meanRecall@K evaluates the performance on the average of all predicate classes, which is more significant for long-tail learning such as in the task of SGG. Both metrics are averaged in the F1@K metric as follows: $F1@K=\frac{2\times R@K\times mR@K}{R@K+mR@K}$ which efficiently represents the trade-off for the model performance between head and tail predictions. Object Detection (OD) accuracy is measured using standard mAP@50 (mAP50).
Implementation details. The latest YOLO versions (V8, V9, and V10) can be used interchangeably in our approach. For simplicity here, we will present experiments for the YOLOV8 medium version [15] (YOLOV8m). YOLOV8m has been trained for object detection for 100 epochs on all datasets with a batch size of 16 and image size 640x640, following the official implementation [^2] https://github.com/ultralytics/ultralytics. All relation prediction models have been trained for 20 epochs, batch size 8 with the SGD optimizer, and an initial learning rate of $1e^{-3}$. For Dynamic Candidate Selection (DCS), we choose $\epsilon=1e^{-5}$. Inference without DCS is performed with a fixed number of 100 proposals for all models. In contrast to YOLO, models trained with Faster-RCNN use 600x1000 image sizes. We fully re-trained seven different relation prediction models on the PSG and IndoorVG datasets: Neural-Motifs [44], VCTree [34], Transformer [35], GPS-Net [26], PE-NET [46], RelTR [4] and EGTR [9] following provided codebases. All models are implemented without any re-weighting [37] or de-biasing strategy [35]. Latency is benchmarked with a batch size of 1333Hardware used: 11th Gen IntelTM CoreTM i9-11950H @ 2.60GHz x 16, NVIDIA GeForce RTX 3080 Laptop GPU 16GB VRAM, 2 x 16GB 3200 MHz RAM..
[…]
Table 4: REACT compared to previous work on VG150 [41, 17] (reported results). mR@K and R@K are the average for $k=[50,100]$ as some approaches are not reporting $k=20$.
B Model mR@K R@K F1@K mAP50
DTS REACT 12.9 27.4 17.6 31.8
TS PE-NET 13.3 32.6 18.9 29.2
SQUAT [16] 15.3 26.7 19.4 -
BGNN [19] 11.6 33.4 17.3 29.0
VCTree 7.2 33.9 11.8 28.1
OS EGTR [10] 7.8 29.3 12.4 30.8
SGTR [20] 13.6 26.5 18.0 25.4
SS-RCNN [37] 9.5 36.0 15.0 23.8
RelTR [4] 8.0 24.3 12.0 26.4
Relation Prediction Accuracy. On the PSG dataset, we experienced an average improvement of 58.47% in F1@k by using our DTS method compared to the same approaches with Faster-RCNN. In addition, our REACT model outperforms the PE-NET model with DTS by a slow margin (23.9 versus 23.2 in F1@K). We observe that the gain in F1@k is not very strongly correlated to the gain in mAP between the Two-Stage and DTS approaches, as we measured a correlation coefficient of 0.75 between the two. On the IndoorVG dataset, we observe a relative improvement of 38.76% in F1@K with the DTS approach. Our REACT model also outperforms PE-NET (DTS) by a strong margin (22.8 versus 20.7 in F1@K). In Table 4 we compare the performance of the REACT model on the VG150 dataset. We observe that REACT outperforms some Two-Stage approaches (Motifs, VCTree, and BGNN) but is not competitive with more recent approaches (PE-NET and SQUAT). We believe this is due to noisy data as well as the poor performance in OD of the YOLO backbone.
Latency and Parameter Count. Regarding latency, we demonstrated an improvement of 84.99% using YOLOV8 instead of Faster-RCNN, which is a considerable gap. Regarding the number of parameters, the DTS approach uses on average 77.47% less parameters than the baseline. Our introduced REACT model is the fastest one with a latency of 32.5ms. By applying our DCS strategy, we gain an average of 66.5% latency on all models (compared to the YOLOV8m-DTS versions) with a relative average loss of 1% in F1@K. The EGTR model stays the best overall in terms of parameter count by a slow margin (42.5M versus 43.3M for REACT). By comparing REACT to the fastest model benchmarked, RelTR [4], REACT is 2.7x faster with $+10$ points on average in F1@K and $+24.9$ points on average in mAP50 for the 3 datasets. By taking into account OD accuracy, RelPred accuracy, and latency, our REACT model is the best compromise. Even for the VG150 dataset, REACT is a better choice because no other two-stage approaches are close to real-time inference. On the other hand, when evaluated fairly, one-stage approaches are clearly behind in terms of mAP and F1@K on all datasets, and are then not a good choice in comparison to REACT.
6 Ablation Studies
Latency and Number of Proposals. As explained before, the second stage of SGG takes as input a fixed number of $n$ proposals as node candidates. There could be a maximum of $n*(n-1)$ possible pairs in the graph, thus when doing matching the computational complexity is supposed to scale accordingly. To demonstrate this hypothesis, we evaluated the performance of the REACT model in both latency and accuracy for different numbers of input proposals $k=\{10,150\}$ with DTS-YOLOV8 and with classical Faster-RCNN. For all experiments, we ranked proposals by confidence of the backbone and selected the top $k$.
Figure 3: Top: Latency for the REACT model using a different number of proposals per image, with batch size 1. Bottom: Average F1@k for the REACT model with DTS-YOLO.
The results of these experiments are shown in Fig. 3 (top), where we observe a plateau phenomenon for REACT with Faster-RCNN above 100 proposals. This behavior does not appear when using DTS+YOLOV8. By using 150 proposals with DTS+YOLOV8, the latency was similar to 10 proposals with Faster-RCNN. The evaluation of the REACT model under different sampling conditions with our DCS strategy is represented in Fig. 3 (bottom). In this figure, we can see that the optimal value was found at 23.89 for $k=42$, just before that the performance starts to be saturated.
Figure 4: Average Recall@K, meanRecall@K, and mAP@50 performance for the REACT model with DCS against the corresponding latency using different variants of the YOLOV8 model.
Scaling Down YOLOV8. To better evaluate the impact of the object detector accuracy on the relation prediction stage, we ran a set of experiments using different variants of YOLOV8 with the REACT model for Scene Graph Generation. Here, we trained the model with the same hyperparameters as before, only modifying the input size of the visual features used for relation prediction to match the different output channel sizes for each YOLOV8 variant. The results are displayed in Figure 4. We observed a strong correlation between mAP@50 and R@K (correlation coefficient = 0.97), and a correlation between mAP@50 and mR@K (correlation coefficient = 0.86). The performance of the larger model, YOLOV8-L is slightly worse in mR@K than the medium YOLOV8-M. We believe that the higher number of dimensions in the features extracted from the backbone could lead to more difficulties for the relation prediction training to converge. In addition, we observed a linear increase in latency with the size of the model. The performance of REACT with the smaller model, YOLOV8-N, is equivalent to the original PE-NET model on the PSG dataset but 22x faster and with 32x fewer parameters (426.5M versus 13.25M). In terms of the overall performance, the medium version seems to be the best. It has a good trade-off between all metrics and latency, making it a good choice for high performance in edge settings.
Features Fusion. During context learning of the REACT model, and more generally in a lot of approaches in SGG [35, 44, 26, 46], different feature sources are used to model the edge representation. First, union and spatial features are combined and then merged with the addition of the subject and object node features. The node features are created using visual and textual features obtained for each proposal. In our REACT model, we removed the dependency on the union features (see Figure 2(d) (i)). In the following, we justify our choice by analyzing the impact of each feature type on learning. The results from these experiments are displayed in Table 5. The baseline REACT model is the second row. With union features (first row) we observe a significant decrease in F1@K as well as latency. We hypothesize that union features are too noisy and confuse the model. This is consistent with previous work which highlights the importance of spatial features for the task [12]. In the last two rows, we removed all dependencies on contextual features (union and spatial) and kept only the concatenation of the subject and object node as initial edge features. We observe only a slight decrease by keeping only text features (last row). In fact, the model can retain up to 71% of the optimal performance by learning only from the textual GLoVe features. This is an interesting finding which questions the balance between text and visual features in the task.
Table 5: Feature sources experiments. V and T are visual and textual features for $<subject,object>$, and S and U are Spatial and Union features, respectively.
T V S U mR@K R@K F1@K Latency Params
14.6 30.8 19.8 36.0 42.7M
19.7 30.2 23.9 32.5 42.7M
14.2 29.9 19.3 27.3 42.7M
14.5 30.1 19.6 18.5 36.6M
13.5 29.1 18.4 16.2 35.2M
Feature Size. Previous two-stage approaches such as PE-NET use a fixed size of 4096 for both the edge and node representation. However, this design was not justified in previous work. Here, we ran experiments to benchmark the efficiency of different feature sizes for learning. We trained the REACT model with four different sizes, see Table 6. We can see that a higher feature size does not necessarily imply better performance. In fact, the best performance in mR@K and F1@K is reported by the smallest size 512. However, the latency and number of parameters scale according to the size, with a maximum of 281.7 ms for the biggest one 4096.
Table 6: Performance of the REACT model using different feature sizes for the edge and node representations.
Size mR@K R@K F1@K Latency Params
512 19.7 30.2 23.9 32.5 42.7M
1024 14.6 30.6 19.7 46.6 71.7M
2048 15.5 29.4 20.3 99.2 168.5M
4096 14.1 30.8 19.3 281.7 515.9M
7 Conclusion
In this work, we proposed a novel implementation of a real-time Scene Graph Generation model that attains competitive performance with a latency of 23ms. We proposed a new architecture, the Decoupled Two-Stage SGG (DTS), aiming at taking advantage of real-time object detectors for SGG. We also proposed a new inference method, Dynamic Candidate Selection (DCS) to further reduce latency by 66.5% on average without compromising performance. Finally, we proposed a new model, REACT, which combines the DTS architecture with efficient prototype learning for lower latency and improved accuracy. Our main finding is that the performance and latency of SGG models are heavily correlated with the quality and quantity of object proposals generated by the object detector. However, regarding both overall performance and latency, a trade-off can be found by using our DCS strategy. Another finding is the over-reliance of models on the text features for relation prediction. Future work will consider implementing the REACT model in constraint settings such as embodied agent navigation or reasoning. Due to its small size, REACT can be embedded onboard robotic platforms and provide reliable and fast predictions that can foster reasoning.
Acknowledgments
This work has been supported by the Brittany Region.
Supplementary Material
Appendix A Datasets
A.1 VG150 Annotations Biases
The VG150 dataset has become the main benchmark for Scene Graph Generation. This dataset has been extracted from original Visual Genome (VG) annotations. In their work, Xu et al. [41] proposed to select the top 150 objects and top 50 predicate classes from VG to compose the VG150 dataset. However, in this process, no manual refinement has been performed to verify the coherence of the data. The original Visual Genome annotations are known to be very noisy [45] with a lot of interchangeable classes. In Tab. 7 we display some of the classes of the VG150 dataset. For every column, we can see that all classes from subsequent rows can be used interchangeably with the label from the first row. This leads to strong confusion of (1) the object detector and (2) the relation predictor.
Objects Predicates
person animal vehicle wears on
girl dog car wearing on top of
woman bear truck above
man cat train
Table 7: Confused classes of VG150.
In Figure 5 we display the confusion matrix for the YOLOV8-medium model trained on VG150. We can clearly observe confusion between the class ”man”, ”boy” and ”person” or another confusion between ”airplane” and ”plane” which demonstrates the low quality of annotations. As a result, we decided not to compare our approach to the VG150 dataset. We still report the results of the REACT model on the VG150 dataset to compare with approaches that cannot be retrained on more qualitative datasets.
Figure 5: Normalized confusion matrix for the YOLOV8-m model on VG150. For convenience, we only display values above 0.1.
A.2 Other Datasets
In addition, we choose not to compare our approach on the GQA [8] and OpenImage [18] datasets. The rationale is as follows: the GQA dataset is also a split of Visual Genome and thus contains the same biases as VG150 [42]. In addition, GQA contains 301 predicate classes with a lot of them being ambiguous or that can be used interchangeably (such as ”standing” and ”standing on”, etc…). The OpenImage dataset contains only very sparse annotations (an average of 2.7 relations per image) which is not tailored for SGG. In addition, it includes a lot of noisy object classes with only a few samples each.
VG150, OpenImage, and GQA are all heavily long-tailed datasets. Our work does not focus on long-tail learning [35], in consequence, we report our main results on the PSG and IndoorVG datasets which are less biased toward tail classes. In addition, these datasets have more qualitative annotations with no spurious or ambiguous classes. For the IndoorVG dataset, we used the V4 version provided in the authors’ codebase. In Table 8 you can find a comparison of each dataset.
Table 8: Comparison of different Scene Graph datasets, # Rel. displays the total number of relationships and Rel/Img the average number of relations per image.
Dataset Images Obj. Pred. # Rel. Rel/Img
VG150 [17] 108,073 150 50 622,705 5.76
OpenImage [18] 133,503 601 30 367,914 2.75
GQA [8] 85,638 200 310 471,614 5.51
PSG [42] 48,749 133 56 278,356 5.71
IndoorVG [28] 14,674 83 34 112,804 7.69
Appendix B YOLO Versions Comparison
Table 9: Performance of the REACT model using different object detectors with the DTS architecture. We used the medium version of each detector.
Model mAP50 mR@K R@K F1@K Lat. Params
YOLOV8 53.1 19.7 30.2 23.9 32.5 42.7M
YOLOV9 53.3 16.3 29.8 21.0 46.5 39.7M
YOLOV10 54.5 16.9 33.0 22.3 55.3 33.4M
In the following, we compare the performance of our REACT model with different YOLO models for feature extraction and object detection. Similar to YOLOV8, we trained the YOLOV9-medium [39] and YOLOV10-medium [38] models on the PSG dataset [42] with batch size 8 for 100 epochs. Then, the REACT model has been trained for 20 epochs using the same hyperparameters as for YOLOV8. Results are displayed in Table 9. We observe no significant improvements in mR@K or R@K by using YOLOV9 or V10 in comparison to V8. The performance of YOLOV9 and YOLOV10 in object detection is superior to YOLOV8, by 0.2 and 1.4 points in mAP, respectively. Interestingly, the overall latency is worse for YOLOV10 and YOLOV9 than by using YOLOV8, which contradicts findings by the respective authors [39, 38]. For YOLOV9, we believe this could come from the size of the feature maps extracted from the backbone which is different than YOLOV8 (YOLOV9m uses a GELAN architecture with an output size of 240, whereas YOLOV8 output size is 192). Higher feature sizes will lead to more computation in the following stages as the subject and object visual representation will be based on these features.
Appendix C Detailed Latency
In the following, we detail the latency of each component of the Two-Stage and Decoupled Two-Stage architectures. Before feeding images to the model, some pre-processing is required (mainly image Transforms). To generate the final predictions, post-process is applied to convert class labels to names and rescale bounding boxes to the original image size. In Table 10 we display these values as well as a detail of the latency of the object detector (feature extraction + object detection) and the relation predictor. In addition, we display the minimum and maximum values obtained on the PSG dataset test set. Very low values account for no bounding box detected and high values for the maximum of 100 bounding boxes detected.
We observe significant differences in latency for the relation prediction stage while object detection stays relatively similar for all models. The difference between max and min values for DTS is 7.91x whereas for TS models it is 12.67x. In addition to a more lightweight backbone, this difference demonstrates that our DTS method is able to remove computation from the relation prediction head. The significant range between minimum and maximum latency is an important parameter that needs to be accounted for deployment. Our REACT model has the lowest difference between min and max values, with 2.85x, showing that our relation head is the most efficient for real-time constraints.
Table 10: Details of the latency for each model. The Min and Max correspond to the min and max values of the full forward pass of the model (without pre and post-process).
B Model Latency (ms)
Pre-process Obj. Det. Rel. Pred. Post-process Min Max
DTS REACT 0.60 16.08 15.70 0.53 19.40 55.36
PE-NET 0.60 17.18 81.42 0.78 25.14 199.54
GPS-NET 0.64 15.31 22.06 0.57 17.17 105.30
Neural-Motifs 0.60 14.82 19.68 0.57 14.00 115.01
VCTree 0.60 16.20 416.47 0.61 63.89 591.94
Transformer 0.57 14.94 17.31 0.51 17.33 138.74
TS PE-NET 0.97 91.35 287.63 0.62 118.09 1625.90
GPS-NET 0.95 86.10 223.75 0.55 105.83 1325.97
Neural-Motifs 1.01 86.19 224.96 0.53 113.87 1373.58
VCTree 0.99 82.10 320.77 0.51 127.54 1583.71
Transformer 0.94 85.20 201.00 0.53 105.96 1308.07
Figure 6: Comparison of Two-Stage and Decoupled Two-Stage architectures with our DCS strategy for inference.
Appendix D Dynamic Candidate Selection
D.1 Rationale
In the Object Detection task, a confidence threshold is traditionally selected at inference time to control the quality of output bounding boxes. The paradigm of SGG leverages an Object Detection component, however, applying a set threshold to control the number of predicted proposals before the relation prediction stage is not desirable. First, the Two-Stage approaches rely on a separate classifier to decode the box labels after the relation prediction stage. The performance of this classifier will impact the final score of the predicted proposals. Second, setting a hard threshold for object detection in our DTS approach will significantly affect the performance concerning relation prediction as the metrics used are not based on precision but rather recall (R@K and mR@K). This will then be unfair when compared with the two-stage approaches. As a result, controlling the number of input proposals by using a saturation threshold for all models (i.e. our DCS strategy) seems more promising. During inference in the wild, a threshold can be applied to the relation scores to only output high-confidence relations and corresponding box coordinates.
D.2 Models Comparison
In this section, we present an in-depth analysis of the DCS strategy. In addition to Decoupled Two-Stage models, we ran evaluation on Two-Stage architectures with Faster-RCNN and DCS on the PSG dataset, results are displayed in Figure 6. We plot the optimum value found for each model. We observe that the optimum for Two-Stage models is significantly higher than for Decoupled Two-Stage models. The average optimum for TS is 74, while for DTS it is 37. This difference can be explained by the low quality of the bounding box predicted by the Faster-RCNN in Two-Stage architectures.
Appendix E Training Details
Table 11: Hyperparameters used for the training of Two-Stage and Decoupled Two-Stage models.
Parameter Value
Optimizer SGD
Learning Rate 0.01
Momentum 0.9
Batch Size 8
Seed 42
Epochs 20
All models are trained using torch.use_deterministic_algorithms as well as a set seed for all random variables. In Table 11 we display the hyperparameters used for all training of Two-Stage and Decoupled Two-Stage models. These values are taken from previous work [34, 44, 46, 35, 26]. We trained the models for 20 epochs using early stopping (grace period=3). For each model, we evaluate the best epoch on the validation set (metric used: F1@K).

### v3

@@
 However, current methods tend to focus on one of the following: (1) improving relation prediction accuracy, (2) enhancing object detection accuracy, or (3) reducing latency, without aiming to balance all three objectives simultaneously.
-To address this limitation, we propose a novel architecture, inference method, and relation prediction model.
-Our proposed solution, the REACT model, achieves the highest inference speed among existing SGG models, improving object detection accuracy without sacrificing relation prediction performance.
-Compared to state-of-the-art approaches, REACT is 2.7 times faster (with a latency of 23 ms) and improves object detection accuracy by 58.51%.
-Furthermore, our proposal significantly reduces model size, with an average of 5.5x fewer parameters [^1] Code is available at https://github.com/Maelic/SGG-Benchmark.
+To address this limitation, we propose the Real-time Efficiency and Accuracy Compromise for Tradeoffs in Scene Graph Generation (REACT) architecture, which achieves the highest inference speed among existing SGG models, improving object detection accuracy without sacrificing relation prediction performance.
+Compared to state-of-the-art approaches, REACT is 2.7 times faster and improves object detection accuracy by 58%.
+Furthermore, our proposal significantly reduces model size, with an average of 5.5x fewer parameters.
+The code is available at https://github.com/Maelic/SGG-Benchmark.
+$\dagger$ Corresponding authors.
+Figure 1: Performance comparison between our approach and previous work, in terms of performance (F1@K metric) and Latency (ms) on the PSG dataset [Yang et al.(2022a)Yang, Ang, Guo, Zhou, Zhang, and Liu].
+Points are scaled by the number of parameters (smaller is better).
 1 Introduction
-Figure 1: Performance comparison between our approach and previous work, in terms of performance (F1@K metric) and Latency (ms) on the PSG dataset [42].
-The points are scaled by the number of parameters for each model (smaller is better).
 Scene Graph Generation (SGG) is the task of generating a structured graph representation from visual inputs.
 Given an image, visual regions corresponding to objects of interest are extracted, and relations between these objects are predicted in the form of <subject, predicate, object> triplets.
-Aggregated together, these triplets form a directed acyclic graph which can be interpreted as a structured representation of the scene.
-In addition, subject and object nodes are traditionally associated with their coordinates on the image (i.e., bounding boxes).
-In recent years, this task has garnered increasing attention thanks to its utility in various downstream applications, such as Visual Question Answering (VQA) [8, 5, 23], Image Captioning [43, 40, 30], or Reasoning of an embodied agent [7, 22, 28].
-However, we observe a significant unbalance between the progress in the SGG task itself and its adoption as a foundational component for downstream tasks [2].
+The set of these triplets forms a directed acyclic graph which can be interpreted as a structured representation of the scene.
+This task has garnered increasing attention recently thanks to its utility in various downstream applications, such as Visual Question Answering (VQA) [Hudson and Manning(2019), Damodaran et al.(2021)Damodaran, Chakravarthy, Kumar, Umapathy, Mitamura, Nakashima, Garcia, and Chu, Liang et al.(2021)Liang, Jiang, and Liu], Image Captioning [Yang et al.(2022b)Yang, Liu, and Wang, Wang et al.(2019)Wang, Beck, and Cohn, Nguyen et al.(2021)Nguyen, Tripathi, Du, Guha, and Nguyen], or Reasoning of an embodied agent [Gadre et al.(2022)Gadre, Ehsani, Song, and Mottaghi, Li et al.(2022b)Li, Guo, Liu, and Sun, Neau et al.(2023a)Neau, Santos, Bosser, and Buche].
+However, we observe a significant imbalance between the progress in the SGG task itself and its adoption in downstream tasks [Chang et al.(2023)Chang, Ren, Xu, Li, Chen, and Hauptmann].
 One of the reasons is the lack of low-cost approaches that could power applications with real-time constraints.
 Because SGG is a task that combines Object Detection (OD) and Relation Prediction (RelPred), a good trade-off between accuracy in these two parameters is also required.
-Nonetheless, current approaches focus on one or two of these parameters without considering the general problem.
-In this work, we explore the bottlenecks in SGG to find an optimal combination of OD accuracy, RelPred accuracy, and latency (see Figure 1).
-We can separate SGG methods into two categories: Two-Stage approaches (TS) and One-Stage approaches (OS).
+In this work, we explore the bottlenecks in SGG to find an optimal combination of OD accuracy, RelPred accuracy, and latency.
+SGG methods can be either Two-Stage (TS) or One-Stage (OS) approaches.
 The former uses a combination of an object detector and a relation predictor to model objects and relations in a sequence.
-The latter uses a single-stage pipeline, to infer relations and object proposals directly from the image features.
-As a result, one-stage approaches are often more efficient for real-time processing [10].
-However, a few challenges remain, especially concerning the accuracy of such approaches in the OD task [20, 9].
-In contrast, two-stage approaches train the object detector and relation predictor sequentially, leading to improved performance in object detection.
-In this work, we propose to investigate the bottlenecks of the two-stage architecture, in terms of OD accuracy and latency.
-First, structural dependencies between blocks of the architecture are reviewed as a Decoupled Two-Stage SGG (DTS-SGG).
-This design permits to maintain performance of the pre-trained detector and allows better control of the information flow between the two stages of the architecture.
-DTS enables the selection of a smaller set of high-quality objects (i.e., proposals) before passing them to the relation prediction stage.
-This approach reduces overall latency, as the complexity of relation prediction increases exponentially with the number of input objects.
+The latter uses a single-stage pipeline to infer relations and object proposals directly from the image features.
+As a result, one-stage approaches are often more efficient for real-time processing [Im et al.(2024b)Im, Nam, Park, Lee, and Park].
+However, they fall short in accuracy in the OD task [Li et al.(2022a)Li, Zhang, and He, Im et al.(2024a)Im, Nam, Park, Lee, and Park].
+In this work, we investigate the bottlenecks of the two-stage architecture in terms of OD accuracy and latency.
+The structural dependencies between blocks of the architecture are reviewed as a Decoupled Two-Stage SGG (DTS-SGG), which allows maintaining the performance of the pre-trained detector, providing better control of the information flow (i.e., detected object proposals) between the two stages of the architecture.
+This approach reduces the overall latency, as the complexity of relation prediction increases exponentially with the number of input objects.
 To select an optimal amount of proposals at inference time, we define a new algorithm, the Dynamic Candidate Selection (DCS).
 DCS uses evaluation results to find the best trade-off between latency and accuracy at test time.
-Thanks to the DTS architecture, we can switch the dependency on a Faster-RCNN model for any real-time object detectors, such as one of the YOLO family (YOLOV8/9/10 [15, 39, 38]).
-Lastly, popular architectures in SGG [46] use a combination of different feature sources, such as union pair features [44] or textual features of the subject and object classes [27].
+Thanks to the DTS architecture, we can switch the dependency on a Faster-RCNN model for any real-time object detectors, such as one of the YOLO family (YOLOV8/9/10 [Jocher et al.(2023)Jocher, Chaurasia, and Qiu, Wang et al.(2024b)Wang, Yeh, and Liao, Wang et al.(2024a)Wang, Chen, Liu, Chen, Lin, Han, and Ding]).
+Recent state-of-the-art SGG models such as PE-NET [Zheng et al.(2023)Zheng, Lyu, Gao, Dai, and Song] use a combination of different feature sources, such as union pair features [Zellers et al.(2018)Zellers, Yatskar, Thomson, and Choi] or textual features of the subject and object classes [Lu et al.(2016)Lu, Krishna, Bernstein, and Fei-Fei].
 This results in a significant amount of processing to gather and refine each type of feature altogether.
-By removing the dependence on sub-optimal features, we can reduce the parameter count of the model and lower the latency.
-By coupling this strategy to the DTS architecture, we propose a new model for SGG, the Real-Time Efficiency and Accuracy Compromise for Tradeoffs in Scene Graph Generation (REACT).
-The contributions of this work are three-fold:
-  1. 1.
-    A new architecture for SGG, the Decoupled Two-Stage (DTS).
-Thanks to this new architecture, real-time object detectors can be coupled with state-of-the-art two-stage RelPred components, and lead to a gain of latency of up to 10x compared to baseline models.
-  2. 2.
-    A new inference method, Dynamic Candidate Selection (DCS), which finds an optimal number of proposals to select for the relation prediction stage, to further improve latency without compromising accuracy.
-  3. 3.
-    A new model, REACT, that extends previous work and our DTS architecture. REACT achieves reduced latency to just 23ms (44 FPS) while enhancing performance on standard metrics (F1@K).
-Our findings demonstrate the inefficiency of current implementations of SGG models and even highlight a few drawbacks in the design of prior approaches, which can be taken into account for future work.
-We also want to emphasize that our method can considerably reduce the time and resources needed for training, which is a step toward more sustainable SGG implementations [33].
+By removing the dependency on sub-optimal features, we can reduce the parameter count of the model and lower the latency.
+By coupling this strategy to the DTS architecture, we propose a new model for SGG, the Real-Time Efficiency and Accuracy Compromise for Tradeoffs in SGG (REACT).
+Thus, the main contributions of this work are: (i) a new architecture for SGG, the Decoupled Two-Stage (DTS); (ii) a new inference method, Dynamic Candidate Selection (DCS), which finds an optimal number of proposals to select for the relation prediction stage, to further improve latency without compromising accuracy; (iii) a new model, REACT, that extends previous work and our DTS architecture, reducing the latency while enhancing performance on standard metrics, see Figure˜1.
 2 Related Work
-Two-stage approaches rely on a Faster-RCNN detector with a ResNeXt-101 backbone for object detection and feature extraction [41].
-Classical approaches include Neural-Motifs [44], VCTree [34], Transformer [35], GPS-Net [26] or PE-NET [46].
-Faster-RCNN is not optimal for real-time constraints because of the size and complexity of the backbone.
-In recent work, Jin et al.
-[11] introduced an approach for real-time SGG based on contextual information.
-In contrast to other approaches, this work does not use visual features to infer relations.
-Instead, it uses the object’s bounding box coordinates predicted by a YOLOV5 [14] model to learn the correlation between pairs of coordinates and predicate classes.
-However, the reliance on spatial features only will neglect fine-grained relations such as the difference between $eating$ or $drinking$.
-Another work [13] proposed to achieve real-time SGG by leveraging a Relation-aware YOLO structure (RYOLO).
-The approach involves predicting the coordinates of relations by adding new anchors on the feature maps of YOLO.
-Then, matching is done with spatially proximate objects detected.
-However, such an approach leads to poor performance in RelPred, with only 3.1% in the mR@100 metric.
-This can be due to the absence of contextual information from the subject and object during relation prediction [44].
+Two-stage SGG approaches typically use Faster-RCNN with a VGG-16 [Xu et al.(2017)Xu, Zhu, Choy, and Fei-Fei, Yang et al.(2018)Yang, Lu, Lee, Batra, and Parikh] or ResNeXt-101 backbone for object detection and feature extraction, as seen in Neural-Motifs [Zellers et al.(2018)Zellers, Yatskar, Thomson, and Choi], VCTree [Tang et al.(2019)Tang, Zhang, Wu, Luo, and Liu], Transformer [Tang et al.(2020)Tang, Niu, Huang, Shi, and Zhang], GPS-Net [Lin et al.(2020)Lin, Ding, Zeng, and Tao], and PE-NET [Zheng et al.(2023)Zheng, Lyu, Gao, Dai, and Song].
+In these approaches, object classes probabilities given by Faster-RCNN are refined during relationship learning, via a dedicated classifier [Zellers et al.(2018)Zellers, Yatskar, Thomson, and Choi, Lin et al.(2020)Lin, Ding, Zeng, and Tao] or by aggregating features from neighboring regions [Yang et al.(2018)Yang, Lu, Lee, Batra, and Parikh].
+However, this design can hinder the performance of SGG models on the task of Object Detection.
+In our work, we choose to effectively Decoupled the Object Detector and Relationship Predictor by removing the dedicated classifier and keeping original class probabilities.
+In addition, Faster-RCNN’s large backbone limits real-time use.
+Recent work [Jin et al.(2023a)Jin, Guo, Meng, Zhu, Xi, Wang, Mu, and Song] proposed a real-time SGG method using contextual cues, relying on YOLOv5 [Jocher(2020)] bounding boxes to learn correlations between box coordinates and predicates—excluding visual features.
+However, the method’s reliance on spatial features will neglect fine-grained relations such as the difference between eatingeating or drinkingdrinking.
+Another approach [Jin et al.(2023c)Jin, Wang, Zhu, Xi, Meng, Mu, and Song] uses a Relation-aware YOLO (RYOLO) to predict relation coordinates via new anchors, matched to nearby objects.
+Yet, lacking subject-object context leads to poor RelPred performance (3.1% mR@100) [Zellers et al.(2018)Zellers, Yatskar, Thomson, and Choi].
 In contrast, we propose to use the features extracted by a YOLO model to generate representations for the subject and object nodes, making the first stage of the architecture more efficient without compromising performance on the second stage.
-Recently, the paradigm of SGG has been tackled as a one-stage architecture, learning to represent both objects and relations at the same time.
-For instance, Sparse-RCNN [36] employs a strategy of triplet querying to generate object proposals and relations.
-Once visual features have been extracted from a CNN backbone, boxes and relations are decoded altogether using a cascade-RCNN scheme.
-RelationTransformer (RelTr) [3] and SGTR [21] are two approaches based on the DETR architecture [1].
-Both approaches generate a sparse set of relations between $<subject,object>$ pairs, in contrast to traditional two-stage approaches which will compute probabilities for all possible relations (dense predictions).
-Inspired by sparse approaches in one-stage SGG, we aim to reduce the computational bottleneck of two-stage SGG by selecting only relevant object pairs for the relation prediction stage.
-Lately, Im et al.
-introduced EGTR [10], a new one-stage approach based on a Deformable-DETR architecture, more efficient than classical DETR for the OD task, paired with a graph relation matching process for RelPred.
-Authors of RelTr and EGTR have compared their work with two-stage approaches such as Motifs [44] and VCTree [34] in terms of latency.
-The comparison shows a clear disadvantage for any two-stage approaches.
-This may be biased by the use of Faster-RCNN for OD in two-stage approaches.
-We strongly believe that lower latency and better accuracy can be attained by modifying two-stage approaches to incorporate real-time object detectors such as YOLO in place of Faster-RCNN.
+One-stage SGG approaches jointly learn object and relation representations.
+Sparse-RCNN [Teng and Wang(2022a)] uses triplet queries to produce objects and relations, decoded via cascade-RCNN after CNN feature extraction.
+RelTr [Cong et al.(2022a)Cong, Yang, and Rosenhahn] and SGTR [Li et al.(2024)Li, Zhang, and He], built on DETR [Carion et al.(2020)Carion, Massa, Synnaeve, Usunier, Kirillov, and Zagoruyko], generate sparse relation sets, unlike dense two-stage predictions.
+More recently, EGTR [Im et al.(2024b)Im, Nam, Park, Lee, and Park] builds upon Deformable-DETR for efficient object detection, paired with graph-based relation matching.
+RelTr and EGTR report lower latency than two-stage models like Motifs [Zellers et al.(2018)Zellers, Yatskar, Thomson, and Choi] and VCTree [Tang et al.(2019)Tang, Zhang, Wu, Luo, and Liu], though these comparisons often rely on Faster-RCNN.
+As highlighted in recent work [Li et al.(2024)Li, Zhang, and He], learning both the object and relation representations at the same time can make the performance of one-stage models in OD sub-optimal in comparison to two-stage approaches.
+We argue that replacing Faster-RCNN with YOLO in two-stage pipelines can achieve both lower latency than two-stage approaches and higher OD accuracy than one-stage approaches, as we shall see in this paper.
 (a) Stage 1 - Two-Stage Faster-RCNN
@@
 (d) Stage 2 - Efficient Prototype Learning (ours)
-Figure 2: Top: the PE-NET architecture [46].
+Figure 2: Top: the PE-NET architecture [Zheng et al.(2023)Zheng, Lyu, Gao, Dai, and Song].
 Bottom: our proposed REACT model with (left) Decoupled Two-Stage (DTS) + YOLO for efficient object detection and feature extraction ; (right) our proposed Efficient Prototype Learning model.
@@
 3 Investigating Bottlenecks in SGG
-In Figure 2 (top row) we display as current the PE-NET model [46]: a state-of-the-art two-stage SGG approach.
-The feature extraction (stage 1 Figure 2(a)) as well as the graph prediction (module 4, Figure 2(b)) is common to all two-stage architectures [34, 35, 6, 19, 44, 26].
-In Stage 1, visual features are extracted by a ResNeXt-101 backbone (Figure 2(a) (a)), and a Feature Pyramid Network (FPN) [25] is used to refine the features at different scales (Figure 2(a) (b)).
-Then, a Region Proposal Network (RPN) [32] generates proposals (Figure 2(a) (c)) which are used to align three different types of features with corresponding image regions: union features, spatial features, and visual features (Figure 2(a) (d), (e), and (f)).
-Union and spatial features are aggregated for every possible pair of proposals and will serve as the initial edge representation in Stage 2 (right) which is the relation prediction stage (Figure 2(b)).
-A classifier extracts the class labels of each proposal (Figure 2(b) (h)) and retrieves corresponding linguistic features from pre-trained models (GloVe word embeddings [31]).
-Then, the visual features are upsampled (Figure 2(b) (i)) and combined with the textual features to form the node representation through semantic fusion (Figure 2(b) (j)).
-Then, prototype learning takes place (Figure 2(b) (k)).
-Finally, object labels and predicate labels are decoded to form the final graph representation (Figure 2(b) (l), (m), and (n)).
-This architecture is complex and may not be efficient for retaining the OD performance of the backbone detector and providing low latency for real-time applications.
-In the following, we present in more detail the bottlenecks related to the Object Detection Accuracy, Proposal Candidate Selection, and the latency of the model.
-3.1 Object Detection Accuracy
-As we can see in Figure 2(b), the strategy employed in the two-stage SGG is to decode object labels twice during the relation prediction stage: one time for extracting text features (Figure 2(b) (h)) and one time to compute the final predictions (Figure 2(b) (m)).
-The predictions obtained from this last step depend on each model’s context learning strategy [34, 44, 26, 35].
-This architecture design is (1) redundant as similar operations are performed twice and (2) can introduce significant differences in object detection accuracy in contrast to the original Faster-RCNN implementation.
-In Table 1, we draw a comparison of the performance of the Faster-RCNN model equipped with a ResNeXt-101 feature extractor on the PSG dataset with different relation prediction heads.
-We can observe that the performance in OD drops significantly, from 1 point for Motifs-TDE [44, 35] to almost 13 points in mAP@50 for GPS-NET [26] after the relation prediction stage.
-Table 1: Performance on Object Detection of Faster-RCNN before and after the relation prediction stage with different SGG models on the PSG dataset [42].
-Backbone Relation Head mAP@50
-Faster-RCNN - 36.4
-Neural-Motifs [44] 35.9
-PE-NET [46] 35.4
-Transformer [35] 34.6
-VC-Tree [34] 32.58
-GPS-NET [26] 31.55
-These results confirm a true dependency between both stages of the pipeline, which seems to hinder the performance of models in the OD task.
-Using the Faster-RCNN detector for feature extraction and object detection may not be efficient for low latency and high accuracy.
-In fact, real-time detectors such as YOLOV8 are today largely outperforming Faster-RCNN [15].
-However, simply replacing Faster-RCNN with YOLO in the architecture of Two-Stage SGG is not possible because YOLO does not possess an RPN or ROI Align component.
-3.2 Proposal Candidate Selection
-Subsequently, performing Non-Maximum Suppression (NMS) (Figure 2(b) (n)) after the context learning is inefficient because a lot of false positive or irrelevant proposals will be selected as node candidates during features refinement and context learning.
-To alleviate the lack of NMS before stage 1, a fixed set of the top proposals (i.e.
-80 or 100) are traditionally considered valid node candidates for stage 2.
+Figure˜2 (top row) is a representation of the PE-NET model [Zheng et al.(2023)Zheng, Lyu, Gao, Dai, and Song]: a state-of-the-art two-stage SGG approach.
+The feature extraction (stage 1 Figure˜2(a)) as well as the graph prediction (module 4, Figure˜2(b)) are common to all two-stage architectures [Tang et al.(2019)Tang, Zhang, Wu, Luo, and Liu, Tang et al.(2020)Tang, Niu, Huang, Shi, and Zhang, Dong et al.(2022)Dong, Gan, Song, Wu, Cheng, and Nie, Li et al.(2021)Li, Zhang, Wan, and He, Zellers et al.(2018)Zellers, Yatskar, Thomson, and Choi, Lin et al.(2020)Lin, Ding, Zeng, and Tao].
+In Stage 1, visual features are extracted by a ResNeXt-101 backbone (Figure˜2(a) (a)), and a Feature Pyramid Network (FPN) [Lin et al.(2017)Lin, Dollár, Girshick, He, Hariharan, and Belongie] is used to refine the features at different scales (Figure˜2(a) (b)).
+Then, a Region Proposal Network (RPN) [Ren et al.(2015)Ren, He, Girshick, and Sun] generates proposals (Figure˜2(a) (c)) which are used to align three different types of features with corresponding image regions: union features, spatial features, and visual features (Figure˜2(a) (d), (e), and (f)) using ROI Align [Ren et al.(2015)Ren, He, Girshick, and Sun].
+Union and spatial features are aggregated for every possible pair of proposals and will serve as the initial edge representation to the relation prediction stage (Stage 2, Figure˜2(b)).
+A classifier extracts the class labels of each proposal (Figure˜2(b) (h)) and retrieves corresponding linguistic features from pre-trained text embeddings (GloVe [Pennington et al.(2014)Pennington, Socher, and Manning]).
+Then, the visual features are combined with the textual features to form the node representation through semantic fusion (Figure˜2(b) (i), (j)).
+Then, prototype learning takes place (Figure˜2(b) (k)).
+Finally, object and predicate labels are decoded to form the final graph representation (Figure˜2(b) (l), (m), and (n)).
+Below, we discuss the main bottlenecks of this architecture for Latency and Object Detection accuracy.
+Object Detection Accuracy.
+The strategy employed in the two-stage SGG is to decode object labels twice during the relation prediction stage: one time for extracting text features (Figure˜2(b) (h)) and one time to compute the final predictions (Figure˜2(b) (m)).
+The predictions obtained from this last step depend on each model’s context learning strategy [Tang et al.(2019)Tang, Zhang, Wu, Luo, and Liu, Zellers et al.(2018)Zellers, Yatskar, Thomson, and Choi, Lin et al.(2020)Lin, Ding, Zeng, and Tao, Tang et al.(2020)Tang, Niu, Huang, Shi, and Zhang].
+This architecture design is (i) redundant as similar operations are performed twice, and (ii) can introduce significant differences in object detection accuracy in contrast to the original Faster-RCNN implementation.
+Table˜1 draws a comparison of the performance of the Faster-RCNN model equipped with a ResNeXt-101 feature extractor on the PSG dataset with different relation prediction heads.
+We can observe that the performance in OD drops significantly, from 1 point wrt Motifs-TDE [Zellers et al.(2018)Zellers, Yatskar, Thomson, and Choi, Tang et al.(2020)Tang, Niu, Huang, Shi, and Zhang] to almost 13 points in mAP@50 wrt GPS-NET [Lin et al.(2020)Lin, Ding, Zeng, and Tao] after the relation prediction stage.
+These results demonstrate a clear dependency between the two stages of the pipeline, which negatively impacts model performance on the object detection (OD) task.
+A second consideration given these results is the low performance of Faster-RCNN.
+Recent state-of-the-art detectors such as YOLOV8/9/10 may be more appropriate for the task.
+In contrast to Faster-RCNN, the YOLO architecture is not designed for Feature Extraction.
+This is the main reason why the Faster-RCNN architecture is still widely used for SGG.
+Table 1: OD accuracy for two-stage models compared to Faster-RCNN, PSG dataset [Yang et al.(2022a)Yang, Ang, Guo, Zhou, Zhang, and Liu].
+Proposal Candidate Selection.
+In the TS architecture, Non-Maximum Suppression (NMS) for object proposals (Figure˜2(b) (n)) is performed in stage 2.
+To alleviate the lack of NMS in stage 1, a fixed set of the top proposals is traditionally considered valid node candidates for stage 2.
 The idea here is that by sampling a large number of proposals per image, it is easier to find valid pairs and relations.
-However, this considerably increases the computational complexity required to generate the graph, which is a bottleneck for real-time applications.
-Having a large number of proposals as candidates can make the final graph less informative, as some important objects can be left out if they do not possess at least one high confidence relation.
-To alleviate this issue, prior works [44, 35, 19] compute the final ranking of relations with the following score formula:
-$\theta_{rel}=\theta_{obj}*\theta_{pred}*\theta_{subj}$ (1)
-[…]
-3.3 Latency of the PE-NET model
-Lastly, we investigate the bottlenecks in latency for the relation prediction stage (see Figure 2(b)) of the PE-NET model [46].
-In PE-NET, node prototypes are formed using a combination of linguistic features and visual features for each independent proposal (see Figure 2(b) (j)).
-Edge features are computed using the concatenation of the spatial and union features (see Figure 2(b) (g)).
+However, this increases the computational complexity of the graph learning.
+Also, the number of top proposals to select (traditionally 80 or 100) is set arbitrarily.
+Recent work [Zellers et al.(2018)Zellers, Yatskar, Thomson, and Choi, Tang et al.(2020)Tang, Niu, Huang, Shi, and Zhang, Li et al.(2021)Li, Zhang, Wan, and He] computes the final ranking of relations with the following score formula: $\theta_{rel}=\theta_{obj}*\theta_{pred}*\theta_{subj},$ where $\theta_{pred}$ is the confidence score of a predicate, given <subject,object><subject,object> pair as candidate; and, $\theta_{obj},\theta_{subj}$ are the respective confidence score of the object detector.
+This formulation assigns greater weight to the confidence scores of the subject and object compared to the predicate.
+As a result, many low-confidence proposals can be filtered out prior to the relation prediction stage with minimal impact on the model’s overall accuracy.
+Latency of the PE-NET model.
+In PE-NET, node prototypes are formed using a combination of linguistic features and visual features (see Figure˜2(b) (j)).
+Edge features are computed using the concatenation of the spatial and union features (see Figure˜2(b) (g)).
 During prototype learning, prototype representations of the subject and object are combined and added to the edge features to form the relation prototype.
 Finally, a gate mechanism is employed to select relevant features from the multi-modal fusion.
-The use of both union and spatial features for the edge representation represents a consequent amount of overload for the model.
-In Figure 2(a) we see that another step of ROI Align is performed on the union of subject and object boxes in step (d).
-However, a lot of visual information is already present in the object and subject features (see Figure 2(a) (f)) which will be fused with edge features during prototype learning.
-In the original work [46], this design is not justified.
-In addition, during edge and node feature refinement, all features are traditionally upsampled to a fixed size of 4096 (see Figure 2(b) (g) and Figure 2(b) (i)).
+The use of both union and spatial features represents a considerable amount of overload for the model.
+In fact, another step of ROI Align is performed on the union of subject and object boxes (Figure˜2(a) step (d)).
+However, visual information is already present in the object and subject features (see Figure˜2(a) (f)), which will be fused with edge features during prototype learning.
+In addition, during edge and node feature refinement, all features are traditionally upsampled to a fixed size of 4096 (see Figure˜2(b) (g) and Figure˜2(b) (i)).
 However, larger feature sizes can significantly augment the latency of the model.
@@
 4 Toward Real-Time SGG
-Based on observations in Section 3, we propose a novel architecture for real-time SGG, which is depicted in Figure 2(c) and Figure 2(d).
-We apply the Decoupled Two-Stage architecture and replace the Faster-RCNN detector of PE-NET with a YOLO backbone, as shown in Figure 2(c).
-We improve the Prototype Learning of PE-NET by removing superfluous components, which results in Efficient Prototype Learning.
-We name the combination of DTS and Efficient Prototype Learning as the REACT model.
-We also propose Dynamic Candidate Selection (DCS) to further reduce the latency of the REACT model at inference time.
-In the following, we detail these components.
-4.1 Decoupled Two-Stage SGG
+This work proposes a novel architecture for real-time SGG, as shown in Figure˜2(c) and Figure˜2(d).
+This architecture applies the Decoupled Two-Stage method, replacing the Faster-RCNN detector of PE-NET with a YOLO backbone (cf.
+Figure˜2(c)).
+The Prototype Learning of PE-NET is improved by removing superfluous components, which results in Efficient Prototype Learning.
+This combination of DTS and Efficient Prototype Learning constitutes the REACT model.
+We also propose Dynamic Candidate Selection (DCS) to further reduce the latency of the proposed model at inference time, as described below.
+Decoupled Two-Stage SGG.
 To maintain the object detector accuracy during relation prediction training, we propose to freeze the regression and the classification head of the object detector and perform NMS before the relation prediction stage.
 With this strategy, the object classifier in the relation prediction stage is also removed and class prediction from the object detector is used for the final prediction.
-The objective of the relation prediction stage becomes then to predict only the pairs and associate predicates, in contrast to also predicting the class labels of objects.
-The original loss used during relation prediction: $\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{rel}}+\mathcal{L}_{\text{obj}}$ becomes: $\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{rel}}$.
-This strategy significantly lowers the complexity and thus the computational load of the model.
-In addition, this approach helps to fairly evaluate the context learning of SGG models, independently of the object detector used.
-With our Decoupled Two-Stage architecture, any object detector can be used in place of Faster-RCNN, without impacting stage 2 of the architecture.
-For instance, one could consider replacing Faster-RCNN with a real-time object detector such as YOLO [11].
-Recent versions of YOLO, such as YOLOV8 [15], YOLOV9 [39] or YOLOV10 [38] are very efficient for real-time object detection.
-To take advantage of YOLO versions, we leveraged the introduced Decoupled Two-Stage strategy and replaced the ResNeXt-101 backbone with YOLO’s CSPNet-based backbones [15, 39] (see Figure 2(c) (a)).
-P5-FPN of Faster-RCNN was also replaced by the YOLOV8/9/10 PAN-FPN Necks (see Figure 2(c) (b)).
-Finally, we removed the RPN and detection head of Faster-RCNN and replaced them with the YOLOV8/9/10 Detection Head (see Figure 2(c) (c)).
+Thus, the relation prediction stage outputs only the pairs and associate predicates, in contrast to also predicting the class labels of objects.
+The loss used during relation prediction: $\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{rel}}+\mathcal{L}_{\text{obj}}$, then becomes: $\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{rel}}$, lowering the model complexity and, consequently, its computational load.
+This approach helps to evaluate the context learning of SGG models independently of the object detector used.
+With this architecture, any object detector can be used in place of Faster-RCNN, without impacting stage 2 of the architecture.
+For instance, one could consider replacing Faster-RCNN with a real-time object detector such as YOLO [Jin et al.(2023a)Jin, Guo, Meng, Zhu, Xi, Wang, Mu, and Song].
+To take advantage of YOLO versions, we leveraged the introduced Decoupled Two-Stage strategy and replaced the ResNeXt-101 backbone with YOLO’s CSPNet-based backbones [Jocher et al.(2023)Jocher, Chaurasia, and Qiu, Wang et al.(2024b)Wang, Yeh, and Liao] (Figure˜2(c) (a)).
+P5-FPN of Faster-RCNN was also replaced by the YOLOV8/9/10 PAN-FPN Necks (Figure˜2(c) (b)).
+Finally, we removed the RPN and detection head of Faster-RCNN and replaced them with the YOLOV8/9/10 Detection Head (Figure˜2(c) (c)).
 Unlike Faster-RCNN, YOLO-based architectures are not designed for feature extraction because they do not implement any ROI Align module.
-To address this issue, we combined the feature maps extracted from the P3, P4, and P5 layers of YOLO and added one 3x3 Conv block to reduce the feature sizes before ROI Align (see Figure 2(c) (f)).
-The ROI Align module from Faster-RCNN has been modified to align features with corresponding bounding boxes at only three different scales (instead of five in the baseline implementation, see Figure 2(c) (f)).
-Thanks to our DTS strategy, switching from one YOLO version to another does not impact the SGG architecture.
-4.2 Dynamic Candidate Selection
-Here, we introduce the Dynamic Candidate Selection (DCS) method to select an optimal number of proposal candidates at inference time.
+To address this issue, we combined the feature maps extracted from the P3, P4, and P5 layers of YOLO and added one 3x3 Conv block to reduce the feature sizes before ROI Align (Figure˜2(c) (f)).
+The ROI Align module from Faster-RCNN has been modified to align features with corresponding bounding boxes at only three different scales (instead of five in the baseline implementation, Figure˜2(c) (f)).
+Dynamic Candidate Selection.
+Here, we introduce the DCS method to select an optimal number of proposal candidates at inference time.
 During training, the relation prediction stage takes as input a maximum amount of proposals $\theta$ ($\theta=100$).
 During evaluation, we uniformly sample the number of proposals $k$, starting from 0.
-Using the generated slope for each metric, we can compute the optimal Dynamic Candidate Selection threshold as follows:
-$x_{\text{opt}}=\min\left\{x\mid\left|f^{\prime}(x)\right|<\epsilon\right\}$ (2)
-During inference, $x_{\text{opt}}$ is applied to control the maximum number of proposals to use as input for the relation prediction stage, successfully lowering the complexity of the task.
-4.3 Efficient Prototype Learning
-In PE-NET, node, and edge features are upsampled 8x and 4x respectively to a fixed size of 4096 by a feed-forward layer (see steps (g) and (h) in Figure 2(b)).
+Using the generated slope for each metric, the optimal DCS threshold is computed as : $x_{\text{opt}}=\min\left\{x\mid\left|f^{\prime}(x)\right|<\epsilon\right\}.$ During inference, $x_{\text{opt}}$ is applied to control the maximum number of proposals to use as input for the relation prediction stage, successfully lowering the complexity of the task.
+Efficient Prototype Learning.
+In PE-NET, node, and edge features are upsampled 8x and 4x respectively to a fixed size of 4096 by a feed-forward layer (steps (g) and (h) in Figure˜2(b)).
 We believe this is unnecessary and that the model will perform similarly without upsampling.
@@
 For the PE-NET model, the use of union features for the edge representation seems redundant with the visual features of the subject and object.
-As a result, we removed the use of union features and dedicated ROI Align step (see Figure 2(a) (d)).
-By doing so, the edge representation refinement step (see Figure 2(a) (g)) can also be removed and the spatial features fed directly to the context learning (see Figure 2(d) (i)).
+As a result, we removed the use of union features and dedicated ROI Align step (Figure˜2(a) (d)).
+By doing so, the edge representation refinement step (Figure˜2(a) (g)) can also be removed and the spatial features fed directly to the context learning (Figure˜2(d) (i)).
 To further boost accuracy, we proposed to modify the fusion of subject $s$ and object $o$ features during Prototype Learning of the PE-NET model.
 We replaced the fusion operation: $ReLU(s+o)-(s-o)^{2}$ with a single linear layer and removed the subsequent gate operation.
-Table 2: Full results of our experiments with the YOLOV8 and Faster-RCNN models on the test set of the PSG dataset.
+5 Experiments
+B D Relation Head mR@20/50/100 R@20/50/100 F1@K mAP50 Latency Params
+CSPDarkNet-53
+YOLOV8m-DTS
+REACT (ours) $\dagger$ 18.3 / 20.0 / 20.8 27.5 / 30.9 / 32.2 23.9 53.1 23.0 43.3M
+REACT (ours) 18.3 / 20.1 / 20.9 27.6 / 30.9 / 32.3 23.9 53.1 32.5 43.3M
+PE-NET $\dagger$ [Zheng et al.(2023)Zheng, Lyu, Gao, Dai, and Song] 17.1 / 19.0 / 19.8 28.0 / 31.3 / 32.9 23.2 $\uparrow$9.1 53.1 $\uparrow$17.6 51.7 $\downarrow$338 187M
+PE-NET [Zheng et al.(2023)Zheng, Lyu, Gao, Dai, and Song] 17.1 / 19.0 / 19.9 28.0 / 31.3 / 33.0 23.2 $\uparrow$9.1 53.1 $\uparrow$17.6 103.0 $\downarrow$261 187M
+GPS-NET $\dagger$ [Lin et al.(2020)Lin, Ding, Zeng, and Tao] 10.9 / 12.6 / 13.5 26.3 / 30.0 / 31.8 17.3 $\uparrow$5.7 53.1 $\uparrow$21.5 27.8 $\downarrow$286 45.3M
+GPS-NET [Lin et al.(2020)Lin, Ding, Zeng, and Tao] 11.1 / 12.7 / 13.7 26.6 / 30.1 / 31.9 17.6 $\uparrow$5.9 53.1 $\uparrow$21.5 41.1 $\downarrow$272 45.3M
+Neural-Motifs $\dagger$ [Zellers et al.(2018)Zellers, Yatskar, Thomson, and Choi] 10.1 / 11.6 / 12.5 25.7 / 28.7 / 30.3 16.2 $\uparrow$3.8 53.1 $\uparrow$17.2 22.4 $\downarrow$296 52.9M
+Neural-Motifs [Zellers et al.(2018)Zellers, Yatskar, Thomson, and Choi] 10.1 / 11.6 / 12.5 25.7 / 28.7 / 30.4 16.2 $\uparrow$3.8 53.1 $\uparrow$17.2 35.8 $\downarrow$282 52.9M
+VCTree $\dagger$ [Tang et al.(2019)Tang, Zhang, Wu, Luo, and Liu] 10.7 / 12.2 / 12.9 22.1 / 25.3 / 26.8 16.1 $\uparrow$3.7 53.1 $\uparrow$20.5 183.2 $\downarrow$336 268.0M
+VCTree [Tang et al.(2019)Tang, Zhang, Wu, Luo, and Liu] 10.7 / 12.2 / 13.0 22.1 / 25.3 / 26.8 16.1 $\uparrow$3.8 53.1 $\uparrow$20.5 239.5 $\downarrow$280 268.0M
+Transformer $\dagger$ [Tang et al.(2020)Tang, Niu, Huang, Shi, and Zhang] 11.3 / 12.4 / 12.8 24.1 / 27.0 / 28.3 16.6 $\uparrow$5.5 53.1 $\uparrow$18.5 21.1 $\downarrow$275 97.9M
+Transformer [Tang et al.(2020)Tang, Niu, Huang, Shi, and Zhang] 11.3 / 12.4 / 13.2 24.1 / 27.1 / 28.9 16.9 $\uparrow$5.7 53.1 $\uparrow$18.5 35.1 $\downarrow$263 97.9M
+ResNeXt-101
+Faster-RCNN
+PE-NET [Zheng et al.(2023)Zheng, Lyu, Gao, Dai, and Song] 10.7 / 11.7 / 12.0 16.8 / 18.7 / 19.5 14.1 35.5 390.4 426.5M
+GPS-NET [Lin et al.(2020)Lin, Ding, Zeng, and Tao] 8.2 / 8.8 / 9.2 15.7 / 18.1 / 19.2 11.7 31.6 313.9 391.6M
+Neural-Motifs [Zellers et al.(2018)Zellers, Yatskar, Thomson, and Choi] 8.9 / 9.6 / 9.8 16.9 / 18.7 / 19.5 12.5 35.9 318.6 369.6M
+VCTree [Tang et al.(2019)Tang, Zhang, Wu, Luo, and Liu] 7.9 / 9.1 / 9.6 18.0 / 20.7 / 22.3 12.4 32.6 519.5 365.9M
+Transformer [Tang et al.(2020)Tang, Niu, Huang, Shi, and Zhang] 7.6 / 8.2 / 8.5 16.2 / 17.9 / 18.7 11.1 34.6 296.2 328.3M
+DETR
+EGTR [Im et al.(2024b)Im, Nam, Park, Lee, and Park] 12.0 / 14.5 / 16.6 24.9 / 30.3 / 33.8 19.4 33.6 78.3 42.5M
+RelTR [Cong et al.(2022b)Cong, Yang, and Rosenhahn] 7.9 / 10.0 / 11.7 12.9 / 17.3 / 20.2 12.4 25.7 59.6 63.7M
+Table 2: Results on the PSG dataset.
 $\dagger$ represents evaluation with our DCS strategy.
 Bold and underlined represent the best and second best for each metric.
-$\downarrow$ and $\uparrow$ represent the improvements of DTS-YOLO against corresponding Two-Stage baselines with Faster-RCNN.
-See Figure 1 for a visual representation.
-B D Relation Head mR@20/50/100 R@20/50/100 F1@K mAP50 Latency Params
-CSPDarkNet-53 YOLOV8m-DTS REACT (ours) $\dagger$  18.3 /  20.1 / 20.8 27.5 / 30.9 / 32.2 23.9 53.1 23.0 43.3M
-REACT (ours) 18.3 / 20.0 / 20.9 27.6 / 30.9 / 32.3 23.9 53.1 32.5 43.3M
-PE-NET $\dagger$ [46] 17.1 / 19.0 / 19.8 28.0 / 31.3 / 32.9 23.2 $\uparrow$9.1 53.1 $\uparrow$17.6 51.7 $\downarrow$338 187M
-PE-NET [46] 17.1 / 19.0 / 19.9 28.0 / 31.3 / 33.0 23.2 $\uparrow$9.1 53.1 $\uparrow$17.6 103.0 $\downarrow$261 187M
-GPS-NET $\dagger$ [26] 10.9 / 12.6 / 13.5 26.3 / 30.0 / 31.8 17.3 $\uparrow$5.7 53.1 $\uparrow$21.5 27.8 $\downarrow$286 45.3M
-GPS-NET [26] 11.1 / 12.7 / 13.7 26.6 / 30.1 / 31.9 17.6 $\uparrow$5.9 53.1 $\uparrow$21.5 41.1 $\downarrow$272 45.3M
-Neural-Motifs $\dagger$ [44] 10.1 / 11.6 / 12.5 25.7 / 28.7 / 30.3 16.2 $\uparrow$3.8 53.1 $\uparrow$17.2 22.4 $\downarrow$296 52.9M
-Neural-Motifs [44] 10.1 / 11.6 / 12.5 25.7 / 28.7 / 30.4 16.2 $\uparrow$3.8 53.1 $\uparrow$17.2 35.8 $\downarrow$282 52.9M
-VCTree $\dagger$ [34] 10.7 / 12.2 / 12.9 22.1 / 25.3 / 26.8 16.1 $\uparrow$3.7 53.1 $\uparrow$20.5 183.2 $\downarrow$336 268.0M
-VCTree [34] 10.7 / 12.2 / 13.0 22.1 / 25.3 / 26.8 16.1 $\uparrow$3.8 53.1 $\uparrow$20.5 239.5 $\downarrow$280 268.0M
-Transformer $\dagger$ [35] 11.3 / 12.4 / 12.8 24.1 / 27.0 / 28.3 16.6 $\uparrow$5.5 53.1 $\uparrow$18.5 21.1 $\downarrow$275 97.9M
-Transformer [35] 11.3 / 12.4 / 13.2 24.1 / 27.1 / 28.9 16.9 $\uparrow$5.7 53.1 $\uparrow$18.5 35.1 $\downarrow$263 97.9M
-ResNeXt-101 Faster-RCNN PE-NET [46] 10.7 / 11.7 / 12.0 16.8 / 18.7 / 19.5 14.1 35.5 390.4 426.5M
-GPS-NET [26] 8.2 / 8.8 / 9.2 15.7 / 18.1 / 19.2 11.7 31.6 313.9 391.6M
-Neural-Motifs [44] 8.9 / 9.6 / 9.8 16.9 / 18.7 / 19.5 12.5 35.9 318.6 369.6M
-VCTree [34] 7.9 / 9.1 / 9.6 18.0 / 20.7 / 22.3 12.4 32.6 519.5 365.9M
-Transformer [35] 7.6 / 8.2 / 8.5 16.2 / 17.9 / 18.7 11.1 34.6 296.2 328.3M
-DETR EGTR [10] 12.0 / 14.5 / 16.6 24.9 / 30.3 / 33.8 19.4 33.6 78.3 42.5M
-RelTR [4] 7.9 / 10.0 / 11.7 12.9 / 17.3 / 20.2 12.4 25.7 59.6 63.7M
-5 Experiments
-Table 3: Full results of our experiments on the test set of IndoorVG. mR@K and R@K represent the average for $k=[20,50,100]$. $\dagger$ represents evaluation with our DCS strategy.
-B Relation Head mR@K R@K F1@K mAP50
-DTS REACT (ours) 18.0 30.9 22.8 37.2
-PE-NET 16.0 29.4 20.7 37.2
-GPS-NET 9.9 29.2 14.8 37.2
-Neural-Motifs 11.1 30.5 16.2 37.2
-VCTree 14.9 18.9 16.7 37.2
-Transformer 12.3 28.4 17.2 37.2
-TS PE-NET 13.8 27.1 18.1 25.3
-GPS-NET 6.7 11.9 8.6 14.8
-Neural-Motifs 9.8 15.2 11.9 26.2
-VCTree 11.0 14.7 12.6 25.5
-Transformer 9.9 16.8 12.5 24.3
-OS EGTR [10] 7.1 10.6 8.5 14.7
-RelTR [4] 7.7 11.4 9.2 14.1
+$\downarrow$ and $\uparrow$ represent the improvements of DTS-YOLO against the corresponding baselines.
 Datasets.
-In SGG, the most widely used dataset is Visual Genome [17].
+In SGG, the most widely used dataset is Visual Genome [Krishna et al.(2017)Krishna, Zhu, Groth, Johnson, Hata, Kravitz, Chen, Kalantidis, Li, Shamma, Bernstein, and Fei-Fei].
 The commonly used split of the data is called VG150 and contains annotations from the top 150 object classes and top 50 predicate classes.
-However, VG150 suffers from severe biases [42, 29, 8] and inherent issues such as the presence of ambiguous classes (e.g.
-the classes ”people”, ”men” etc) [42, 24].
-Recently, two new datasets have been proposed with higher-quality annotations: the PSG [42] and IndoorVG [28] datasets.
+However, VG150 suffers from severe biases [Yang et al.(2022a)Yang, Ang, Guo, Zhou, Zhang, and Liu, Neau et al.(2023b)Neau, Santos, Bosser, and Buche, Hudson and Manning(2019)] and inherent issues such as the presence of ambiguous classes (e.g.
+the classes "people", "men" etc) [Yang et al.(2022a)Yang, Ang, Guo, Zhou, Zhang, and Liu, Liang et al.(2019)Liang, Bai, Zhang, Qian, Zhu, and Mei].
+Recently, two new datasets have been proposed with higher-quality annotations: the PSG [Yang et al.(2022a)Yang, Ang, Guo, Zhou, Zhang, and Liu] and IndoorVG [Neau et al.(2023a)Neau, Santos, Bosser, and Buche] datasets.
 Both datasets have been carefully annotated to remove the presence of ambiguous classes or wrong annotations, which makes them a better choice for benchmarking our approach.
 Metrics.
-Following previous work [41, 34], we used the Recall@K (R@K) and meanRecall@K (mR@K) metrics to measure the performance of models.
+Following previous work [Xu et al.(2017)Xu, Zhu, Choy, and Fei-Fei, Tang et al.(2019)Tang, Zhang, Wu, Luo, and Liu], we used the Recall@K (R@K) and meanRecall@K (mR@K) metrics to measure the performance of models.
 For both metrics, we evaluate the recall on the top $K$ ($k=[20,50,100]$) relations predicted, ranked by confidence.
@@
 The latest YOLO versions (V8, V9, and V10) can be used interchangeably in our approach.
-For simplicity here, we will present experiments for the YOLOV8 medium version [15] (YOLOV8m).
-YOLOV8m has been trained for object detection for 100 epochs on all datasets with a batch size of 16 and image size 640x640, following the official implementation [^2] https://github.com/ultralytics/ultralytics.
-All relation prediction models have been trained for 20 epochs, batch size 8 with the SGD optimizer, and an initial learning rate of $1e^{-3}$.
+For simplicity here, we will present experiments for the YOLOV8 medium version [Jocher et al.(2023)Jocher, Chaurasia, and Qiu] (YOLOV8m).
+YOLOV8m has been trained for object detection for 100 epochs on all datasets with a batch size of 16 and image size 640x640.
+Relation prediction models have been trained for 20 epochs, batch size 8 with the SGD optimizer, and an initial learning rate of $1e^{-3}$.
 For Dynamic Candidate Selection (DCS), we choose $\epsilon=1e^{-5}$.
-Inference without DCS is performed with a fixed number of 100 proposals for all models.
+Inference without DCS is performed with 100 proposals for all models.
 In contrast to YOLO, models trained with Faster-RCNN use 600x1000 image sizes.
-We fully re-trained seven different relation prediction models on the PSG and IndoorVG datasets: Neural-Motifs [44], VCTree [34], Transformer [35], GPS-Net [26], PE-NET [46], RelTR [4] and EGTR [9] following provided codebases.
-All models are implemented without any re-weighting [37] or de-biasing strategy [35].
-Latency is benchmarked with a batch size of 1333Hardware used: 11th Gen IntelTM CoreTM i9-11950H @ 2.60GHz x 16, NVIDIA GeForce RTX 3080 Laptop GPU 16GB VRAM, 2 x 16GB 3200 MHz RAM..
-[…]
-Table 4: REACT compared to previous work on VG150 [41, 17] (reported results). mR@K and R@K are the average for $k=[50,100]$ as some approaches are not reporting $k=20$.
+We fully re-trained seven different relation prediction models on the PSG and IndoorVG datasets: Neural-Motifs [Zellers et al.(2018)Zellers, Yatskar, Thomson, and Choi], VCTree [Tang et al.(2019)Tang, Zhang, Wu, Luo, and Liu], Transformer [Tang et al.(2020)Tang, Niu, Huang, Shi, and Zhang], GPS-Net [Lin et al.(2020)Lin, Ding, Zeng, and Tao], PE-NET [Zheng et al.(2023)Zheng, Lyu, Gao, Dai, and Song], RelTR [Cong et al.(2022b)Cong, Yang, and Rosenhahn], and EGTR [Im et al.(2024a)Im, Nam, Park, Lee, and Park] following the provided codebases.
+All models are implemented without any re-weighting [Teng and Wang(2022b)] or de-biasing strategy [Tang et al.(2020)Tang, Niu, Huang, Shi, and Zhang].
+Latency is benchmarked with a batch size of 1111Hardware used: 11th Gen IntelTM CoreTM i9-11950H @ 2.60GHz x 16, NVIDIA GeForce RTX 3080 Laptop GPU 16GB VRAM, 2 x 16GB 3200 MHz RAM..
+Results.
+In Table˜2 and Table˜4 we display the performance of our DTS architecture in OD accuracy compared to previous work.
+The F1@K results from Table˜2 are represented as a graph in Figure˜1.
+In the text, we refer to improvements as mean absolute percentage.
+B Relation Head mR@K R@K F1@K mAP50
+DTS
+REACT (ours) 18.0 30.9 22.8 37.2
+PE-NET 16.0 29.4 20.7 37.2
+GPS-NET 9.9 29.2 14.8 37.2
+Neural-Motifs 11.1 30.5 16.2 37.2
+VCTree 14.9 18.9 16.7 37.2
+Transformer 12.3 28.4 17.2 37.2
+TS
+PE-NET 13.8 27.1 18.1 25.2
+GPS-NET 6.7 11.9 8.6 17.4
+Neural-Motifs 9.8 15.2 11.9 26.2
+VCTree 11.0 14.7 12.6 25.5
+Transformer 9.9 16.8 12.5 24.2
+OS
+EGTR [Im et al.(2024b)Im, Nam, Park, Lee, and Park] 7.1 10.6 8.5 14.7
+RelTR [Cong et al.(2022b)Cong, Yang, and Rosenhahn] 7.7 11.4 9.2 14.1
+Table 3: Results on IndoorVG. mR@K and R@K are average for $k=[20,50,100]$.
 B Model mR@K R@K F1@K mAP50
@@
 TS PE-NET 13.3 32.6 18.9 29.2
-SQUAT [16] 15.3 26.7 19.4 -
-BGNN [19] 11.6 33.4 17.3 29.0
+SQUAT [Jung et al.(2023)Jung, Kim, Kim, and Cho] 15.3 26.7 19.4 -
+BGNN [Li et al.(2021)Li, Zhang, Wan, and He] 11.6 33.4 17.3 29.0
 VCTree 7.2 33.9 11.8 28.1
-OS EGTR [10] 7.8 29.3 12.4 30.8
-SGTR [20] 13.6 26.5 18.0 25.4
-SS-RCNN [37] 9.5 36.0 15.0 23.8
-RelTR [4] 8.0 24.3 12.0 26.4
-Relation Prediction Accuracy.
+OS EGTR [Im et al.(2024b)Im, Nam, Park, Lee, and Park] 7.8 29.3 12.4 30.8
+SGTR [Li et al.(2022a)Li, Zhang, and He] 13.6 26.5 18.0 25.4
+SS-RCNN [Teng and Wang(2022b)] 9.5 36.0 15.0 23.8
+RelTR [Cong et al.(2022b)Cong, Yang, and Rosenhahn] 8.0 24.3 12.0 26.4
+Table 4: REACT compared to previous work on VG150 [Xu et al.(2017)Xu, Zhu, Choy, and Fei-Fei, Krishna et al.(2017)Krishna, Zhu, Groth, Johnson, Hata, Kravitz, Chen, Kalantidis, Li, Shamma, Bernstein, and Fei-Fei] (reported results).
+mR@K and R@K are the average for $k=[50,100]$ as some approaches do not report $k=20$.
+By using YOLOV8, we improved the mAP by 54.37% on the PSG dataset (Table˜2) and 69.92% on IndoorVG (Table˜4), compared to TS approaches with Faster-RCNN.
+Compared to OS approaches (EGTR [Im et al.(2024a)Im, Nam, Park, Lee, and Park] and RelTR [Cong et al.(2022b)Cong, Yang, and Rosenhahn]), mAP is improved by 120% on average, which demonstrates the inefficiency of such approaches for the OD task.
+For a fair comparison, we also present the performance of the REACT model on the VG150 dataset, see Table˜4.
+The performance of REACT in mAP is better by one point than the best approach EGTR.
+We believe that the low quality of the bounding box annotations and the ambiguous classes in VG150 are confusing the YOLOV8 detector, which leads to only a small improvement against Faster-RCNN.
 On the PSG dataset, we experienced an average improvement of 58.47% in F1@k by using our DTS method compared to the same approaches with Faster-RCNN.
-In addition, our REACT model outperforms the PE-NET model with DTS by a slow margin (23.9 versus 23.2 in F1@K).
+In addition, our REACT model outperforms the PE-NET model with DTS by a small margin (23.9 versus 23.2 in F1@K).
 We observe that the gain in F1@k is not very strongly correlated to the gain in mAP between the Two-Stage and DTS approaches, as we measured a correlation coefficient of 0.75 between the two.
-On the IndoorVG dataset, we observe a relative improvement of 38.76% in F1@K with the DTS approach.
+On the IndoorVG dataset, we observe an improvement of 38.76% in F1@K with the DTS approach.
 Our REACT model also outperforms PE-NET (DTS) by a strong margin (22.8 versus 20.7 in F1@K).
-In Table 4 we compare the performance of the REACT model on the VG150 dataset.
-We observe that REACT outperforms some Two-Stage approaches (Motifs, VCTree, and BGNN) but is not competitive with more recent approaches (PE-NET and SQUAT).
-We believe this is due to noisy data as well as the poor performance in OD of the YOLO backbone.
-Latency and Parameter Count.
+We observe that REACT outperforms some Two-Stage approaches (Motifs, VCTree, and BGNN) for RelPred on VG150 (see Table˜4) but is not competitive with more recent approaches (PE-NET and SQUAT).
+We believe this is due to noisy feature representation.
+Because REACT has fewer parameters, it is more sensitive to noisy data.
 Regarding latency, we demonstrated an improvement of 84.99% using YOLOV8 instead of Faster-RCNN, which is a considerable gap.
 Regarding the number of parameters, the DTS approach uses on average 77.47% less parameters than the baseline.
-Our introduced REACT model is the fastest one with a latency of 32.5ms.
+Our introduced REACT model is the fastest one, with a latency of 32.5ms.
 By applying our DCS strategy, we gain an average of 66.5% latency on all models (compared to the YOLOV8m-DTS versions) with a relative average loss of 1% in F1@K.
 The EGTR model stays the best overall in terms of parameter count by a slow margin (42.5M versus 43.3M for REACT).
-By comparing REACT to the fastest model benchmarked, RelTR [4], REACT is 2.7x faster with $+10$ points on average in F1@K and $+24.9$ points on average in mAP50 for the 3 datasets.
+By comparing REACT to the fastest model benchmarked, RelTR [Cong et al.(2022b)Cong, Yang, and Rosenhahn], REACT is 2.7x faster with +10+10 points on average in F1@K and +24.9+24.9 points on average in mAP50 for the 3 datasets.
 By taking into account OD accuracy, RelPred accuracy, and latency, our REACT model is the best compromise.
@@
 On the other hand, when evaluated fairly, one-stage approaches are clearly behind in terms of mAP and F1@K on all datasets, and are then not a good choice in comparison to REACT.
-6 Ablation Studies
 Latency and Number of Proposals.
 As explained before, the second stage of SGG takes as input a fixed number of $n$ proposals as node candidates.
-There could be a maximum of $n*(n-1)$ possible pairs in the graph, thus when doing matching the computational complexity is supposed to scale accordingly.
+There could be a maximum of n(n1)n*(n-1) possible pairs in the graph, thus when doing matching the computational complexity is supposed to scale accordingly.
 To demonstrate this hypothesis, we evaluated the performance of the REACT model in both latency and accuracy for different numbers of input proposals $k=\{10,150\}$ with DTS-YOLOV8 and with classical Faster-RCNN.
 For all experiments, we ranked proposals by confidence of the backbone and selected the top $k$.
-Figure 3: Top: Latency for the REACT model using a different number of proposals per image, with batch size 1. Bottom: Average F1@k for the REACT model with DTS-YOLO.
-The results of these experiments are shown in Fig.
-3 (top), where we observe a plateau phenomenon for REACT with Faster-RCNN above 100 proposals.
+Figure 3: Left: Latency for the REACT model using a different number of proposals per image, with batch size 1. Right: Average F1@k for the REACT model with DTS-YOLO.
+The results of these experiments are shown in fig.˜3 (top), where we observe a plateau phenomenon for REACT with Faster-RCNN above 100 proposals.
 This behavior does not appear when using DTS+YOLOV8.
 By using 150 proposals with DTS+YOLOV8, the latency was similar to 10 proposals with Faster-RCNN.
-The evaluation of the REACT model under different sampling conditions with our DCS strategy is represented in Fig.
-3 (bottom).
+The evaluation of the REACT model under different sampling conditions with our DCS strategy is represented in fig.˜3 (bottom).
 In this figure, we can see that the optimal value was found at 23.89 for $k=42$, just before that the performance starts to be saturated.
-Figure 4: Average Recall@K, meanRecall@K, and mAP@50 performance for the REACT model with DCS against the corresponding latency using different variants of the YOLOV8 model.
-Scaling Down YOLOV8.
-To better evaluate the impact of the object detector accuracy on the relation prediction stage, we ran a set of experiments using different variants of YOLOV8 with the REACT model for Scene Graph Generation.
-Here, we trained the model with the same hyperparameters as before, only modifying the input size of the visual features used for relation prediction to match the different output channel sizes for each YOLOV8 variant.
-The results are displayed in Figure 4.
-We observed a strong correlation between mAP@50 and R@K (correlation coefficient = 0.97), and a correlation between mAP@50 and mR@K (correlation coefficient = 0.86).
-The performance of the larger model, YOLOV8-L is slightly worse in mR@K than the medium YOLOV8-M.
-We believe that the higher number of dimensions in the features extracted from the backbone could lead to more difficulties for the relation prediction training to converge.
-In addition, we observed a linear increase in latency with the size of the model.
-The performance of REACT with the smaller model, YOLOV8-N, is equivalent to the original PE-NET model on the PSG dataset but 22x faster and with 32x fewer parameters (426.5M versus 13.25M).
-In terms of the overall performance, the medium version seems to be the best.
-It has a good trade-off between all metrics and latency, making it a good choice for high performance in edge settings.
 Features Fusion.
-During context learning of the REACT model, and more generally in a lot of approaches in SGG [35, 44, 26, 46], different feature sources are used to model the edge representation.
+During context learning of the REACT model, and more generally in a lot of approaches in SGG [Tang et al.(2020)Tang, Niu, Huang, Shi, and Zhang, Zellers et al.(2018)Zellers, Yatskar, Thomson, and Choi, Lin et al.(2020)Lin, Ding, Zeng, and Tao, Zheng et al.(2023)Zheng, Lyu, Gao, Dai, and Song], different feature sources are used to model the edge representation.
 First, union and spatial features are combined and then merged with the addition of the subject and object node features.
 The node features are created using visual and textual features obtained for each proposal.
-In our REACT model, we removed the dependency on the union features (see Figure 2(d) (i)).
+In our REACT model, we removed the dependency on the union features (see Figure˜2(d) (i)).
 In the following, we justify our choice by analyzing the impact of each feature type on learning.
-The results from these experiments are displayed in Table 5.
+The results from these experiments are displayed in Table˜5.
 The baseline REACT model is the second row.
@@
 We hypothesize that union features are too noisy and confuse the model.
-This is consistent with previous work which highlights the importance of spatial features for the task [12].
+This is consistent with previous work which highlights the importance of spatial features for the task [Jin et al.(2023b)Jin, Guo, Meng, Zhu, Xi, Wang, Mu, and Song].
 In the last two rows, we removed all dependencies on contextual features (union and spatial) and kept only the concatenation of the subject and object node as initial edge features.
@@
 In fact, the model can retain up to 71% of the optimal performance by learning only from the textual GLoVe features.
-This is an interesting finding which questions the balance between text and visual features in the task.
-Table 5: Feature sources experiments. V and T are visual and textual features for $<subject,object>$, and S and U are Spatial and Union features, respectively.
+This suggests a trade off between text and visual features in the task.
 T V S U mR@K R@K F1@K Latency Params
@@
 13.5 29.1 18.4 16.2 35.2M
-Feature Size.
-Previous two-stage approaches such as PE-NET use a fixed size of 4096 for both the edge and node representation.
-However, this design was not justified in previous work.
-Here, we ran experiments to benchmark the efficiency of different feature sizes for learning.
-We trained the REACT model with four different sizes, see Table 6.
-We can see that a higher feature size does not necessarily imply better performance.
-In fact, the best performance in mR@K and F1@K is reported by the smallest size 512.
-However, the latency and number of parameters scale according to the size, with a maximum of 281.7 ms for the biggest one 4096.
-Table 6: Performance of the REACT model using different feature sizes for the edge and node representations.
-Size mR@K R@K F1@K Latency Params
-512 19.7 30.2 23.9 32.5 42.7M
-1024 14.6 30.6 19.7 46.6 71.7M
-2048 15.5 29.4 20.3 99.2 168.5M
-4096 14.1 30.8 19.3 281.7 515.9M
+Table 5: Feature sources experiments. V and T are visual and textual features for <subject,object><subject,object>, and S and U are Spatial and Union features, respectively.
 6 Conclusion
@@
 Acknowledgments
-This work has been supported by the Brittany Region.
-Supplementary Material
-Appendix A Datasets
-A.1 VG150 Annotations Biases
-The VG150 dataset has become the main benchmark for Scene Graph Generation.
-This dataset has been extracted from original Visual Genome (VG) annotations.
-In their work, Xu et al.
-[41] proposed to select the top 150 objects and top 50 predicate classes from VG to compose the VG150 dataset.
-However, in this process, no manual refinement has been performed to verify the coherence of the data.
-The original Visual Genome annotations are known to be very noisy [45] with a lot of interchangeable classes.
-In Tab.
-7 we display some of the classes of the VG150 dataset.
-For every column, we can see that all classes from subsequent rows can be used interchangeably with the label from the first row.
-This leads to strong confusion of (1) the object detector and (2) the relation predictor.
-Objects Predicates
-person animal vehicle wears on
-girl dog car wearing on top of
-woman bear truck above
-man cat train
-Table 7: Confused classes of VG150.
-In Figure 5 we display the confusion matrix for the YOLOV8-medium model trained on VG150.
-We can clearly observe confusion between the class ”man”, ”boy” and ”person” or another confusion between ”airplane” and ”plane” which demonstrates the low quality of annotations.
-As a result, we decided not to compare our approach to the VG150 dataset.
-We still report the results of the REACT model on the VG150 dataset to compare with approaches that cannot be retrained on more qualitative datasets.
-Figure 5: Normalized confusion matrix for the YOLOV8-m model on VG150. For convenience, we only display values above 0.1.
-A.2 Other Datasets
-In addition, we choose not to compare our approach on the GQA [8] and OpenImage [18] datasets.
-The rationale is as follows: the GQA dataset is also a split of Visual Genome and thus contains the same biases as VG150 [42].
-In addition, GQA contains 301 predicate classes with a lot of them being ambiguous or that can be used interchangeably (such as ”standing” and ”standing on”, etc…).
-The OpenImage dataset contains only very sparse annotations (an average of 2.7 relations per image) which is not tailored for SGG.
-In addition, it includes a lot of noisy object classes with only a few samples each.
-VG150, OpenImage, and GQA are all heavily long-tailed datasets.
-Our work does not focus on long-tail learning [35], in consequence, we report our main results on the PSG and IndoorVG datasets which are less biased toward tail classes.
-In addition, these datasets have more qualitative annotations with no spurious or ambiguous classes.
-For the IndoorVG dataset, we used the V4 version provided in the authors’ codebase.
-In Table 8 you can find a comparison of each dataset.
-Table 8: Comparison of different Scene Graph datasets, # Rel. displays the total number of relationships and Rel/Img the average number of relations per image.
-Dataset Images Obj. Pred. # Rel. Rel/Img
-VG150 [17] 108,073 150 50 622,705 5.76
-OpenImage [18] 133,503 601 30 367,914 2.75
-GQA [8] 85,638 200 310 471,614 5.51
-PSG [42] 48,749 133 56 278,356 5.71
-IndoorVG [28] 14,674 83 34 112,804 7.69
-Appendix B YOLO Versions Comparison
-Table 9: Performance of the REACT model using different object detectors with the DTS architecture. We used the medium version of each detector.
-Model mAP50 mR@K R@K F1@K Lat. Params
-YOLOV8 53.1 19.7 30.2 23.9 32.5 42.7M
-YOLOV9 53.3 16.3 29.8 21.0 46.5 39.7M
-YOLOV10 54.5 16.9 33.0 22.3 55.3 33.4M
-In the following, we compare the performance of our REACT model with different YOLO models for feature extraction and object detection.
-Similar to YOLOV8, we trained the YOLOV9-medium [39] and YOLOV10-medium [38] models on the PSG dataset [42] with batch size 8 for 100 epochs.
-Then, the REACT model has been trained for 20 epochs using the same hyperparameters as for YOLOV8.
-Results are displayed in Table 9.
-We observe no significant improvements in mR@K or R@K by using YOLOV9 or V10 in comparison to V8.
-The performance of YOLOV9 and YOLOV10 in object detection is superior to YOLOV8, by 0.2 and 1.4 points in mAP, respectively.
-Interestingly, the overall latency is worse for YOLOV10 and YOLOV9 than by using YOLOV8, which contradicts findings by the respective authors [39, 38].
-For YOLOV9, we believe this could come from the size of the feature maps extracted from the backbone which is different than YOLOV8 (YOLOV9m uses a GELAN architecture with an output size of 240, whereas YOLOV8 output size is 192).
-Higher feature sizes will lead to more computation in the following stages as the subject and object visual representation will be based on these features.
-Appendix C Detailed Latency
-In the following, we detail the latency of each component of the Two-Stage and Decoupled Two-Stage architectures.
-Before feeding images to the model, some pre-processing is required (mainly image Transforms).
-To generate the final predictions, post-process is applied to convert class labels to names and rescale bounding boxes to the original image size.
-In Table 10 we display these values as well as a detail of the latency of the object detector (feature extraction + object detection) and the relation predictor.
-In addition, we display the minimum and maximum values obtained on the PSG dataset test set.
-Very low values account for no bounding box detected and high values for the maximum of 100 bounding boxes detected.
-We observe significant differences in latency for the relation prediction stage while object detection stays relatively similar for all models.
-The difference between max and min values for DTS is 7.91x whereas for TS models it is 12.67x.
-In addition to a more lightweight backbone, this difference demonstrates that our DTS method is able to remove computation from the relation prediction head.
-The significant range between minimum and maximum latency is an important parameter that needs to be accounted for deployment.
-Our REACT model has the lowest difference between min and max values, with 2.85x, showing that our relation head is the most efficient for real-time constraints.
-Table 10: Details of the latency for each model. The Min and Max correspond to the min and max values of the full forward pass of the model (without pre and post-process).
-B Model Latency (ms)
-Pre-process Obj. Det. Rel. Pred. Post-process Min Max
-DTS REACT 0.60 16.08 15.70 0.53 19.40 55.36
-PE-NET 0.60 17.18 81.42 0.78 25.14 199.54
-GPS-NET 0.64 15.31 22.06 0.57 17.17 105.30
-Neural-Motifs 0.60 14.82 19.68 0.57 14.00 115.01
-VCTree 0.60 16.20 416.47 0.61 63.89 591.94
-Transformer 0.57 14.94 17.31 0.51 17.33 138.74
-TS PE-NET 0.97 91.35 287.63 0.62 118.09 1625.90
-GPS-NET 0.95 86.10 223.75 0.55 105.83 1325.97
-Neural-Motifs 1.01 86.19 224.96 0.53 113.87 1373.58
-VCTree 0.99 82.10 320.77 0.51 127.54 1583.71
-Transformer 0.94 85.20 201.00 0.53 105.96 1308.07
-Figure 6: Comparison of Two-Stage and Decoupled Two-Stage architectures with our DCS strategy for inference.
-Appendix D Dynamic Candidate Selection
-D.1 Rationale
-In the Object Detection task, a confidence threshold is traditionally selected at inference time to control the quality of output bounding boxes.
-The paradigm of SGG leverages an Object Detection component, however, applying a set threshold to control the number of predicted proposals before the relation prediction stage is not desirable.
-First, the Two-Stage approaches rely on a separate classifier to decode the box labels after the relation prediction stage.
-The performance of this classifier will impact the final score of the predicted proposals.
-Second, setting a hard threshold for object detection in our DTS approach will significantly affect the performance concerning relation prediction as the metrics used are not based on precision but rather recall (R@K and mR@K).
-This will then be unfair when compared with the two-stage approaches.
-As a result, controlling the number of input proposals by using a saturation threshold for all models (i.e.
-our DCS strategy) seems more promising.
-During inference in the wild, a threshold can be applied to the relation scores to only output high-confidence relations and corresponding box coordinates.
-D.2 Models Comparison
-In this section, we present an in-depth analysis of the DCS strategy.
-In addition to Decoupled Two-Stage models, we ran evaluation on Two-Stage architectures with Faster-RCNN and DCS on the PSG dataset, results are displayed in Figure 6.
-We plot the optimum value found for each model.
-We observe that the optimum for Two-Stage models is significantly higher than for Decoupled Two-Stage models.
-The average optimum for TS is 74, while for DTS it is 37.
-This difference can be explained by the low quality of the bounding box predicted by the Faster-RCNN in Two-Stage architectures.
-Appendix E Training Details
-Table 11: Hyperparameters used for the training of Two-Stage and Decoupled Two-Stage models.
-Parameter Value
-Optimizer SGD
-Learning Rate 0.01
-Momentum 0.9
-Batch Size 8
-Seed 42
-Epochs 20
-All models are trained using torch.use_deterministic_algorithms as well as a set seed for all random variables.
-In Table 11 we display the hyperparameters used for all training of Two-Stage and Decoupled Two-Stage models.
-These values are taken from previous work [34, 44, 46, 35, 26].
-We trained the models for 20 epochs using early stopping (grace period=3).
-For each model, we evaluate the best epoch on the validation set (metric used: F1@K).
+This work has been supported by the Brittany Region and the Knut and Alice Wallenberg foundation and the Wallenberg AI, Autonomous Systems and Software Program (WASP).